### v1

# Convexified Message-Passing Graph Neural Networks

Abstract
Graph Neural Networks (GNNs) have become prominent methods for graph representation learning, demonstrating strong empirical results on diverse graph prediction tasks. In this paper, we introduce Convexified Message-Passing Graph Neural Networks (CGNNs), a novel and general framework that combines the power of message-passing GNNs with the tractability of convex optimization. By mapping their nonlinear filters into a reproducing kernel Hilbert space, CGNNs transform training into a convex optimization problem, which can be solved efficiently and optimally by projected gradient methods. This convexity further allows the statistical properties of CGNNs to be analyzed accurately and rigorously. For two-layer CGNNs, we establish rigorous generalization guarantees, showing convergence to the performance of the optimal GNN. To scale to deeper architectures, we adopt a principled layer-wise training strategy. Experiments on benchmark datasets show that CGNNs significantly exceed the performance of leading GNN models, achieving 10–40% higher accuracy in most cases, underscoring their promise as a powerful and principled method with strong theoretical foundations. In rare cases where improvements are not quantitatively substantial, the convex models either slightly exceed or match the baselines, stressing their robustness and wide applicability. Though over-parameterization is often employed to enhance performance in non-convex models, we show that our CGNNs framework yields shallow convex models that can surpass these models in both accuracy and resource efficiency.
1 Introduction
Graphs serve as versatile models across various domains, such as social networks [47], protein design [22], and medical diagnosis [27]. Graph representation learning has thus attracted widespread attention in recent years, with Graph Neural Networks (GNNs) emerging as powerful tools. Message Passing Neural Networks (MPNNs) [18, 20, 45], a most widely adopted class of GNNs, provide a scalable and elegant architecture that effectively exploits a message-passing mechanism to extract useful information and learn high-quality representations from graph data.
In this paper, we present the new and general framework of Convexified Message-Passing Graph Neural Networks (CGNNs), reducing the training of message-passing GNNs to a convex optimization problem solvable efficiently and optimally via projected gradient methods. The convexity of CGNNs offers a distinct advantage: it ensures global optimality and computational tractability without relying on, e.g., over-parameterization, unlike many existing approaches [1, 10, 11, 60]. This convexity also enables a precise and theoretically grounded analysis of generalization and statistical behavior.
We exhibit our framework by convexifying two-layer Graph Convolution Networks (GCNs) [7, 24], a widely-used GNN model for learning over graph-structured data. Inspired by the convexification of convolutional neural networks by Zhang et al. [59], we adapt their methodology to the graph domain. A high-level overview of our convexification pipeline is summarized in Figure 1, explained as follows. The key challenge lies in handling GCNs’ nonlinearity: we address this by first relaxing the space of GCN filters to a reproducing kernel Hilbert space (RKHS) (step (1) in Figure 1), effectively reducing the model to one with a linear activation function (step (2)). We then show that GCNs with RKHS filters admit a low-rank matrix representation (step (3)). Finally, by relaxing the non-convex rank constraint into a convex nuclear norm constraint (step (4)), we arrive at our CGCN formulation.
Theoretically, we establish an upper bound on the expected generalization error of our class of two-layer CGCNs, composed of the best possible performance attainable by a two-layer GCN with infinite data and a complexity term that decays to zero polynomially with the number of training samples. Our result dictates that the generalization error of a CGCN provably converges to that of an optimal GCN, offering both strong statistical guarantees and practical relevance. Moreover, our analysis naturally extends to deeper models by applying the argument layer-wise, providing insights into the statistical behavior of multi-layer CGCNs as well.
Finally, we perform an extensive empirical analysis over various benchmark datasets, where our framework exhibits outstanding results. First, we apply our framework to a wide range of leading message-passing GNN models, including GAT [45], GATv2 [4], GIN [51], GraphSAGE [20], and ResGatedGCN [3]. We also extended our framework to hybrid transformer models, specifically GraphGPS [36] and GraphTrans [40, 49], which integrate local message passing GNNs with global attention mechanisms. Remarkably, in most cases, the convex variants outperform the original non-convex counterparts by a large magnitude, obtaining 10–40% higher accuracy, highlighting the practical power of our approach alongside its theoretical appeal. Even in the minority of cases where improvements are modest, the convexified models still match or slightly exceed the baseline performance, further validating the robustness and broad applicability of our method. While over-parameterization is commonly used to boost performance in non-convex models [1, 10, 11, 60], we show that our CGNNs framework enables the construction of shallow convex architectures that outperform over-parameterized non-convex ones in both accuracy and resource efficiency.
GNN(1)Map toRKHS(2)Linearactivation(3)Low-rank matrixparameterization(4)Nuclear normrelaxationTrainCGNN
Figure 1: Convexification Procedure of a message-passing GNN into a CGNN.
2 Related Work
Message-Passing GNNs. Early developments of message-passing GNNs include GCN [7, 24], GAT [45], GIN [51], GraphSAGE [20], and others. Such methods effectively leverage a message-passing mechanism to recursively aggregate useful neighbor information [18], enabling them to learn high-quality representations from graph data. Those models have proven effective across diverse tasks, obtaining state-of-the-art performance in various graph prediction tasks (see, e.g., [5, 20, 36, 45, 51, 55]). Our framework of CGNNs harnesses the power of message-passing GNNs by transforming training into a convex optimization problem. This shift not only enables efficient computation of globally optimal solutions, but also allows an accurate and rigorous analysis of statistical properties.
Convergence of Gradient Methods to Global Optima. A vast line of research analyzed the convergence of gradient-based methods to global optima when training neural networks, but under restrictive conditions. Several works rely on over-parameterization, often requiring network widths far exceeding the number of training samples [1, 11, 60], with milder assumptions in shallow networks [10, 43]. However, the convexity of our CGNNs enables global optimality and computational efficiency without relying on any of the above restrictive assumptions.
Convex Neural Networks. The area of convex neural networks is relatively underexplored. Zhang et al. [57] proposed a convex relaxation step for fully-connected neural networks, which inspired the convexification of convolutional neural networks (CNNs) [59]. Other works study linear CNNs [19] and linear GNNs [41], which are of less interest in practice since they collapse all layers into one. Recently, Cohen and Agmon [6] proposed convex versions of specific GNN models termed as aggregation GNN, which apply only to a signal with temporal structure [15]. Their work is narrowly tailored and focused on a specific application in swarm robotics, where convexity arises from modeling assumptions tailored to domain adaptation. In contrast, our CGNN framework establishes a general and theoretically principled convexification of diverse message-passing GNNs by transforming the standard architecture itself, independent of application-specific constraints. Unlike their formulation that does not preserve, e.g., standard GNN modularity, our approach retains these key structural features and applies broadly to a wide range of graph learning tasks. As such, CGNNs represent a scalable and domain-agnostic foundation for convex GNN training, grounded in a fundamentally different mathematical framework.
Graph Transformers. Instead of aggregating local neighborhood, Graph Transformers capture interaction information between any pair of nodes via a single self-attention layer. Some works design specific attention mechanisms or positional encodings [21, 25, 28, 52], while others combine message-passing GNNs to build hybrid architectures [5, 36, 40, 49]. Those methods enable nodes to interact with any other node, facilitating direct modeling of long-range relations. Our framework broadly applies to such hybrid methods, as we exhibit on the leading models GraphGPS [36] and GraphTrans [40, 49]. Empirically, applying our framework consistently improves performance in graph classification tasks, with significant gains on GraphGPS, showcasing the versatility and effectiveness of our approach.
Notations. For brevity, $[n]$ denotes the discrete set $\{1,2,\dots,n\}$ for any $n\in\mathbb{N}_{>0}$. For a rectangular matrix $A$, let $\|A\|_{*}$ be its nuclear norm (i.e., the sum of $\mathcal{A}_{\ell}$’s singular values), and $\|A\|_{2}$ be its spectral norm (i.e., maximal singular value). Let $\ell^{2}(\mathbb{N})$ be the set of countable dimensional of square-summable sequences $v$, i.e., $\sum_{i=1}^{\infty}v_{i}^{2}<\infty$.
3 Preliminaries
Network data can be captured by an underlying undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathcal{W})$, where $\mathcal{V}=\{1,...,n\}$ denotes the vertex set, $\mathcal{E}$ denotes the edge set and $\mathcal{W}:\mathcal{E}\rightarrow\mathbb{R}$ denotes the edge weight function, while a missing edge $(i,j)$ is depicted by $\mathcal{W}(i,j)=0$. The neighborhood of node $i\in\mathcal{V}$ is denoted by $N_{i}:=\{j\in\mathcal{V}|(j,i)\in\mathcal{E}\}$. The data on top of this graph is modeled as a graph signal $\mathcal{X}\in\mathbb{R}^{n\times F}$, whose $i^{\text{th}}$ row $[\mathcal{X}]_{i}=\mathbf{x}_{i}$ is an $F$-dimensional feature vector assigned to node $i\in\mathcal{V}$. We leverage the framework of graph signal processing (GSP) [33] as the mathematical foundation for learning from graph signals. For relating the graph signal $\mathcal{X}$ with the underlying structure of $\mathcal{G}$, we define a graph shift operator (GSO) by $\mathcal{S}\in\mathbb{R}^{n\times n}$, which satisfies $[\mathcal{S}]_{ij}=s_{ij}=0$ if $(j,i)\notin\mathcal{E}$ for $j\neq i$. Commonly utilized GSOs include the adjacency matrix [37, 38], the Laplacian matrix [42], and their normalized counterparts [7, 16]. By viewing $\mathcal{S}$ as a linear operator, the graph signal $\mathcal{X}$ can be shifted over the nodes, yielding that the output at node $i$ for the $f^{\text{th}}$ feature becomes $[\mathcal{S}\mathcal{X}]_{if}=\sum_{j=1}^{n}[\mathcal{S}]_{ij}[\mathcal{X}]_{jf}=\sum_{j\in N_{i}}s_{ij}x_{jf}$. Note that, whereas $\mathcal{S}\mathcal{X}$ corresponds to a local information exchange between a given node and its direct neighbors, applying this linear operator recursively aggregates information from nodes in further hops along the graph. Namely, the $k$-shifted signal $\mathcal{S}^{k}\mathcal{X}$ aggregates information from nodes that are $k$-hops away.
Graph Convolutional Networks (GCNs) [17] stack multiple graph (convolutional) layers to extract high-level node representations from graph signals. To formally describe the creation of convolutional features in a GCN with $L\in\mathbb{N}$ layers, let the output of layer $\ell-1$ (and thus the input to the $\ell^{\text{th}}$ layer) be $\mathcal{X}_{\ell-1}$, comprising of $F_{\ell-1}$ feature vectors. The input to layer 00 is $\mathcal{X}_{0}=\mathcal{X}$, yielding $F_{0}=F$, while the final output has $G:=F_{L}$ features. Given a set of filters $A:=\{\mathcal{A}_{\ell k}\in\mathbb{R}^{F_{\ell-1}\times F_{\ell}}|k\in[K]\cup\{0\},\ell\in[L]\}$, the graph (convolutional) filter at layer $\ell$ processes the $F_{\ell-1}$ features of the input graph signal $\mathcal{X}_{\ell-1}$ in parallel by $F_{\ell}$ graph filters to output the following $F_{\ell}$ convolutional features:
$\Psi^{\mathcal{A}_{\ell}}(\mathcal{X};\mathcal{S}):=\sum_{k=0}^{K}\mathcal{S}^{k}\mathcal{X}_{\ell-1}\mathcal{A}_{\ell k},$ (1)
where $\mathcal{A}_{\ell}:=(\mathcal{A}_{\ell k})_{k=0}^{K}$ concatenates the filters at layer $\ell$. Note that all nodes share the same parameters to weigh equally the information from all $k$-hop away neighbors for any $k\in[K]$, which corresponds to the parameter sharing of GCNs. A GCN is then defined as a nonlinear mapping between graph signals $\Phi:\mathbb{R}^{n\times F}\rightarrow\mathbb{R}^{n\times G}$, which is obtained by feeding the aggregated features through a pointwise nonlinear activation function $\sigma:\mathbb{R}\rightarrow\mathbb{R}$ (which, for brevity, denotes its entrywise application in (2)), producing the feature vectors of the $\ell^{\text{th}}$ layer’s output:
$\Phi(\mathcal{X};\mathcal{S},A):=\mathcal{X}_{L};\quad\mathcal{X}_{\ell}:=\sigma(\Psi^{\mathcal{A}_{\ell}}(\mathcal{X};\mathcal{S})).$ (2)
We denote the $f^{\text{th}}$ column of the filters $\mathcal{A}_{\ell k}$ as $\mathbf{a}^{f}_{\ell k}=[a_{\ell k}^{fg}]_{g\in[F_{\ell-1}]}$. Given some positive constants $\{R_{\ell}\}_{\ell\in[L]}$, we consider the following non-convex function class of (regularized) GCNs:
[…]
By setting $R_{\ell}=\infty$ for any $\ell\in[L]$, this function class reduces to the class of (unregularized) GCNs.
3.1 Training GCNs
We focus on the graph classification task with $G$ classes. Here, we are given a training set $\mathcal{T}:=\{(\mathcal{G}^{(j)},\mathcal{X}^{(j)},y^{(j)})\}_{j=1}^{m}$ of $m$ samples, where the $j^{\text{th}}$ sample consists of a graph $\mathcal{G}^{(j)}$, a graph signal $\mathcal{X}^{(j)}\in\mathbb{R}^{n\times F}$ over $\mathcal{G}^{(j)}$ and the graph’s true label $y^{(j)}\in[G]$. Without loss of generality, we assume that each graph $\mathcal{G}^{(j)}$ has $n$ nodes (otherwise, we can add isolated dummy nodes). We then train a GCN (as formulated in (2)), while node features from the last layer are aggregated to obtain the graph-level representation via a permutation invariant function, such as summation or a more sophisticated graph-level pooling function [54, 56]. Training is done by solving the following optimization problem for some loss function $J:\mathbb{R}^{n\times G}\times\mathbb{R}^{n}\rightarrow\mathbb{R}$:
$\Phi^{*}_{\text{gcn}}\in\operatorname*{arg\,min}_{\Phi\in\mathcal{F}_{\text{gcn}}}\frac{1}{m}\sum_{j=1}^{m}J(\Phi(\mathcal{X}^{(j)};\mathcal{S},A),y^{(j)}).$ (4)
We assume that $J$ is convex and $M$-Lipschitz in its first argument (predictions) for any fixed second argument (labels). The number of parameters to be trained are determined by the number of layers $L$, the number of filters $K$ and features $F_{\ell}$ in each layer $\ell$. By (2), each GCN $\Phi\in\mathcal{F}_{\text{gcn}}$ depends nonlinearly on the parameters, making the optimization problem in (4) non-convex. Next, we thus present a relaxation of the class $\mathcal{F}_{\text{gcn}}$, enabling a convex formulation of the associated problem in (4).
4 Convexification of GCNs
As $\mathcal{F}_{\text{gcn}}$ in (3) is a non-convex set and the activation function is nonlinear, the ERM problem on GCNs is non-convex, raising the challenges mentioned in Section 1. To overcome them, we propose a procedure for making both the ERM’s predictions domain and loss function convex. Following Zhang et al. [59], we develop the class of Convexified GCNs (CGCNs). In Section 4.1, we first depict our method for linear activation functions. Given that the linear activation function would collapse all layers into one, this case is of no interest in practice. However, it supplies us with some insight into the general kernelization scheme of GCNs. In the nonlinear case (Section 4.2), we reduce the problem to the linear case by proposing a relaxation to a suitably chosen Reproducing Kernel Hilbert Space (RKHS), enabling us to obtain a convex formulation of the associated ERM problem. We then present an algorithm for learning CGCNs (Section 4.3), discuss its scalability and time complexity (Section 4.4), and establish its generalization guarantees (Section 4.5).
4.1 Convexificaition of GCNs with Linear Activation Functions
To provide the intuition for our method, we employ the nuclear norm in the simple case of the linear activation function $\sigma(\mathbf{w})=\mathbf{w}$, which provides us with a convex function class, as opposed to (3). In this case, the output of layer $\ell$ is simply $\mathcal{X}_{\ell}=\sum_{k=0}^{K}\mathcal{S}^{k}\mathcal{X}_{\ell-1}\mathcal{A}_{\ell k}$ due to (1)-(2). Hence, $\mathcal{X}_{\ell}$ linearly depends on the filters $\{\mathcal{A}_{\ell k}\}_{k=0}^{K}$ of layer $\ell$, i.e., the convolution function $\Psi^{\mathcal{A}_{\ell}}:\mathbb{R}^{n\times F_{\ell-1}}\rightarrow\mathbb{R}^{n\times F_{\ell}}$ in (1) linearly depends on the parameters.
Low-Rank Constraints. The filters matrix $\mathcal{A}_{\ell k}\in\mathbb{R}^{F_{\ell-1}\times F_{\ell}}$ has rank at most $F_{\ell-1}$. As $\mathcal{A}_{\ell}:=(\mathcal{A}_{\ell k})_{k=0}^{K}$ concatenates the filters at layer $\ell$, then $\operatorname*{rank}(\mathcal{A}_{\ell})\leq F_{\ell-1}$. Hence, the class $\mathcal{F}_{\text{gcn}}$ of GCNs in (3) corresponds to a collection of functions depending on norm and low-rank constraints imposed on the filters. We next define:
$\mathcal{F}_{\text{linear-gcn}}:=\{\Psi^{\mathcal{A}_{\ell}}:$ $\max_{0\leq k\leq K}\|\mathbf{a}^{f}_{\ell k}\|_{2}\leq R_{\ell}\quad\text{ and }$ (5a)
$\operatorname*{rank}(\mathcal{A}_{\ell})=F_{\ell-1}\quad\forall f,\ell\},$ (5b)
which is the specific form of our original class of GCNs in (3) for the linear case. While (5a) holds even when the parameters are not shared over all nodes, the low-rank constraint (5b) will no longer be satisfied in this case. Namely, (5b) is achieved via GCNs’ parameter sharing: each node processes its neighborhood up to $K$ hops away using the same filters, and thus the number of independent rows of parameters remains at most $F_{\ell-1}$, regardless of the number of nodes in the graph.
Nuclear Norm Relaxation. However, filters that satisfy the constraints (5a) and (5b) form a non-convex set. A standard convex relaxation of a rank constraint involves replacing it with a nuclear norm constraint. That is, controlling the sum of singular values instead of the number of non-zero ones, similarly to how the $\ell_{1}$-c can be viewed as a convex relaxation of the $\ell_{0}$-norm constraint.
Upper Bounding the Nuclear-Norm. In Appendix A, we prove that:
Lemma 4.1.
Assume that the filters $\mathcal{A}_{\ell}$ at each layer $\ell$ satisfy (5a) and (5b). Then, $\mathcal{A}_{\ell}$’s nuclear norm is upper bounded as $\|\mathcal{A}_{\ell}\|_{*}\leq\mathcal{B}_{\ell}$ for some $\mathcal{B}_{\ell}>0$ that is a function of $R_{\ell},F_{\ell},F_{\ell-1},K$.
Using Lemma 4.1, we can replace (5a) and (5b) with a nuclear norm constraint, thus forming a convex function class. Namely, we define the class of Convexified GCNs (CGCNs) as:
$\mathcal{F}_{\text{cgcn}}:=\{\{\Psi^{\mathcal{A}_{\ell}}\}_{\ell=1}^{L}:\|\mathcal{A}_{\ell}\|_{*}\leq\mathcal{B}_{\ell}\quad\forall\ell\},$ (6)
while guaranteeing that $\mathcal{F}_{\text{gcn}}\subseteq\mathcal{F}_{\text{cgcn}}$. Thus, solving the ERM in (4) over $\mathcal{F}_{\text{cgcn}}$ instead of $\mathcal{F}_{\text{gcn}}$ yields a convex optimization problem over a wealthier class of functions. Next, for the more general setting of nonlinear activation functions, we devise an algorithm capable of solving this convex program.
4.2 Convexification of GCNs with Nonlinear Activation Functions
For certain nonlinear activation functions $\sigma$ and suitably chosen kernel function $\kappa$ such as the Gaussian RBF kernel, we next prove how the class $\mathcal{F}_{\text{gcn}}$ of GCNs in (3) can be relaxed to a Reproducing Kernel Hilbert Space (RKHS), reducing the problem to the linear activation case (Readers may refer to [39] for a brief overview on RKHS). Beforehand, we rephrase the filter at layer $\ell$ from (1)-(2) as follows. Denoting the $f^{\text{th}}$ column of the filters $\mathcal{A}_{\ell k}$ as $\mathbf{a}^{f}_{\ell k}$, in Appendix C we prove that:
Lemma 4.2.
A GCN induces a nonlinear filter $\tau_{\ell}^{f}$ at each layer $\ell$ for the $f^{\text{th}}$ feature given as follows (where the activation function $\sigma$ is applied entrywise):
$\begin{array}[]{c}\tau_{\ell}^{f}:\mathcal{Z}\in\mathbb{R}^{(K+1)\times F_{\ell-1}}\mapsto\sigma(\sum_{k=0}^{K}\langle[\mathcal{Z}]_{k+1},\mathbf{a}^{f}_{\ell k}\rangle).\end{array}$ (7)
For any graph signal $\mathcal{S}^{k}\mathcal{X}$, let $\mathbf{z}_{i}^{k}(\mathcal{X}):=[\mathcal{S}^{k}\mathcal{X}]_{i}$ be the $i^{\text{th}}$ row of $\mathcal{S}^{k}\mathcal{X}$, which is the information aggregated at node $i$ from $k$-hops away. Defining the matrix $\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1})\in\mathbb{R}^{(K+1)\times F_{\ell-1}}$ whose $(k+1)^{\text{th}}$ row is $[\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1})]_{k+1}=\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1})$ for $k\in[K]\cup\{0\}$, the output at layer $\ell$ is thus given by $\tau_{\ell}^{k}(\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1}))$.
Kernelization of GCNs. In the sequel, we consider a training data $\mathcal{T}=\{(\mathcal{G}^{(j)},\mathcal{X}^{(j)},y^{(j)})\}_{j=1}^{m}$, some layer of index $\ell$ and a positive semi-definite kernel function $\kappa_{\ell}:\mathbb{R}^{F_{\ell-1}}\times\mathbb{R}^{F_{\ell-1}}\rightarrow\mathbb{R}$ for layer $\ell$. Thus, for brevity, we omit indices involving $\ell$ and $\mathcal{X}_{\ell-1}^{(j)}$ when no ambiguity arises. Namely, we denote, e.g., $\mathcal{Z}_{i,j}:=\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1}^{(j)})$ and $\mathbf{z}_{i,j}^{k}:=\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}^{(j)})$.
First, we prove in Lemma 4.3 that the filter specified by (7) resides in the RKHS induced by the kernel (See Appendix D for a proof, which adapts the one by Zhang et al. [59] to GCNs). Zhang et al. [59] prove that convolutional neural networks can be contained within the RKHS induced by properly chosen kernel and activation functions. For certain choices of kernels (e.g., the Gaussian RBF kernel) and a sufficiently smooth $\sigma$ (e.g., smoothed hinge loss), we can similarly prove that the filter $\tau_{\ell}^{f}$ in (7) is contained within the RKHS induced by the kernel function $\kappa_{\ell}$. See Appendices B.1-B.2 for possible choices of kernels and Appendix B.3 for valid activation functions, as well as an elaboration on sufficient smoothness in Remark B.2.
Lemma 4.3.
The filter $\tau^{f}_{\ell}$ at layer $\ell$ for the $f^{\text{th}}$ feature can be reparametrized as a linear combination of kernel products, reducing the problem to the linear activations case. Formally:
$\tau^{f}_{\ell}(\mathcal{Z}_{i,j})=\sum_{k=0}^{K}\sum_{(i^{\prime},j^{\prime})\in[n]\times[m]}\alpha_{k,(i^{\prime},j^{\prime})}^{f}\cdot\kappa_{\ell}(\mathbf{z}_{i,j}^{k},\mathbf{z}_{i^{\prime},j^{\prime}}^{k}),$ (8)
where $\{\alpha_{k,(i^{\prime},j^{\prime})}^{f}\}_{(i^{\prime},j^{\prime},k)\in[n]\times[m]\times([K]\cup\{0\})}$ are some coefficients.
Remark 4.4.
Lemma 4.3 suggests that filters of the form (8) are parameterized by a finite set of coefficients $\{\alpha_{k,(i^{\prime},j^{\prime})}^{f}\}$, allowing optimization to be carried over these coefficients instead of the original parameters $\mathcal{A}_{\ell}$ (See Section 4.3 for details on solving this optimization problem).
To avoid re-deriving all results in the kernelized setting, we reduce the problem to the linear case in Section 4.1. Thus, consider the (symmetric) kernel matrix $\mathcal{K}^{k}_{\ell}\in\mathbb{R}^{nm\times nm}$, whose entry at row $(i,j)$ and column $(i^{\prime},j^{\prime})$ equals to $\kappa_{\ell}(\mathbf{z}_{i,j}^{k},\mathbf{z}_{i^{\prime},j^{\prime}}^{k})$. We factorize the kernel matrix as $\mathcal{K}^{k}_{\ell}=Q_{\ell}^{k}(Q_{\ell}^{k})^{\top}$, where $Q_{\ell}^{k}\in\mathbb{R}^{nm\times P}$ (as in [59], we can use, e.g., Cholesky factorization [8] with $P=nm$).
Remark 4.5.
The $(i,j)$-th row of $Q_{\ell}^{k}$, i.e., $\mathbf{q}_{i,j}^{k}:=[Q_{\ell}^{k}]_{(i,j)}\in\mathbb{R}^{P}$, can be interpreted as a feature vector instead of the original $\mathbf{z}_{i,j}^{k}\in\mathbb{R}^{F_{\ell-1}}$ (the information aggregated at node $i$ from $k$-hops away).
Thus, the reparameterization in (8) of the filter at layer $\ell$ for the $f^{\text{th}}$ feature can be rephrased as:
$\tau^{f}_{\ell}(\mathcal{Z}_{i,j})=\sum_{k=0}^{K}\langle\mathbf{q}_{i,j}^{k},\hat{\mathbf{a}}_{\ell k}^{f}\rangle;\quad\hat{\mathbf{a}}_{\ell k}^{f}:=\sum_{(i^{\prime},j^{\prime},k)\in[n]\times[m]\times([K]\cup\{0\})}\alpha_{k,(i^{\prime},j^{\prime})}^{f}\mathbf{q}_{i,j}^{k}$ (9)
Next, we outline the learning and testing phases of CGCNs (See Appendix D for more details).
Remark 4.6.
Relaxing the filters to the induced RKHS reduces the problem to the linear case. The resulting relaxed filters satisfy a low-rank constraint and have a bounded nuclear norm. By Section 4.1, convexly relaxing the low-rank constraint via the nuclear norm yields a convex function class.
To summarize, we describe the forward pass operation of a single layer in Algorithm 1. It processes an input graph signal $\mathcal{X}$ and transforms it in two steps:
Step 1: Arranging Input to the Layer. Each layer $\ell$ takes as input the graph signal $\mathcal{X}_{\ell-1}$ from the previous layer and computes intermediate graph signals $Z_{\ell}^{k}(\mathcal{X}_{\ell-1})$ for different hop distances $k$ (line 2). For each node $i$, information is aggregated from its $k$-hop neighbors (line 4), from which the $i^{\text{th}}$ row of $Z_{\ell}^{k}(\mathcal{X}_{\ell-1})$ is then obtained based on kernel products (lines 4-6).
Step 2: Computing the Layer Output. Using the intermediate representations and the trainable filters $\hat{\mathcal{A}}_{\ell k}$, the layer’s output signal $\mathcal{X}_{\ell}$ is calculated (line 8). Specifically, each node’s output feature vector at layer $\ell$ is computed via a weighted sum of the transformed signals (line 9).
Algorithm 1 Forward Pass of the $\ell^{\text{th}}$ CGCN Layer
Input: Graph signal $\mathcal{X}\in\mathbb{R}^{n\times F}$, filters $\hat{\mathcal{A}}_{\ell}=(\hat{\mathcal{A}}_{\ell k})_{k=0}^{K}$
1:Step 1: Arrange input to layer $\ell$
2:Given an input graph signal $\mathcal{X}_{\ell-1}\in\mathbb{R}^{n\times F}$ from the previous layer, compute a new graph signal $Z_{\ell}^{k}(\mathcal{X}_{\ell-1})\in\mathbb{R}^{n\times P}$ for any $\ell\in[L]$ and $k\in[K]\cup\{0\}$ as follows.
3:Compute $\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}^{(j)}):=[\mathcal{S}^{k}\mathcal{X}_{\ell-1}^{(j)}]_{i}$, the signal aggregated at node $i$ from $k$-hops away with respect to the training graph signal $\mathcal{X}^{(j)}$.
4:As usual in kernel methods, we need to compute $\kappa_{\ell}(\mathbf{z},\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}))$, where $\mathbf{z}\in\mathbb{R}^{n\times F}$ is some input graph signal.
5:Form the vector $\mathbf{v}_{\ell}^{k}(\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}))$ whose elements are kernel products (See Appendix D).
6:Compute the $i^{\text{th}}$ row of $Z_{\ell}^{k}(\mathcal{X}_{\ell-1})$ using $Z_{\ell}^{k}(\mathcal{X}_{\ell-1})_{i}=(Q^{k})^{\dagger}\mathbf{v}_{\ell}^{k}(\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}))$, where $(Q^{k})^{\dagger}$ is the pseudo-inverse of $Q^{k}$ (See Appendix D for more details).
7:Step 2: Compute the layer’s output
8:Using the computed signals $Z_{\ell}^{k}(\mathcal{X}_{\ell-1})$, construct the output graph signal $\mathcal{X}_{\ell}$ for layer $\ell$.
9:Layer $\ell$ induces a linear mapping between graph signals $\hat{\Psi}^{\hat{\mathcal{A}}_{\ell}}:\mathbb{R}^{n\times F_{\ell-1}}\rightarrow\mathbb{R}^{n\times F_{\ell}}$ s.t. the value of the $f^{\text{th}}$ feature of node $i$ at layer $\ell$ is $[\mathcal{X}_{\ell}]_{if}=[\hat{\Psi}^{\hat{\mathcal{A}}_{\ell}}(\mathcal{X}_{\ell-1})]_{if}=\sum_{k=0}^{K}\langle\hat{\mathbf{z}}_{i}^{k}(\mathcal{X}_{\ell-1}),\hat{\mathbf{a}}_{\ell k}^{f}\rangle$, where $\hat{\mathbf{a}}_{\ell k}^{f}$ are the trainable parameters (filters) of the CGCN layer from (9).
Output: Output graph signal $\mathcal{X}_{\ell}\in\mathbb{R}^{n\times F_{\ell}}$
4.3 The Algorithm for Learning Multi-Layer CGCNs
Algorithm 2 Training of CGCNs
Input: Training set $\mathcal{T}=\{(\mathcal{G}^{(j)},\mathcal{X}^{(j)},y^{(j)})\}_{j=1}^{m}$; Kernels $\{\kappa_{\ell}\}_{\ell=1}^{L}$; Regularization parameters {}$\{\mathcal{B}_{\ell}\}_{\ell=1}^{L}$; Number of graph filters $\{F_{\ell}\}_{\ell=1}^{L}$
1:for each layer $2\leq\ell\leq L$ do
2:     Taking $\{(\mathcal{G}^{(j)},\mathcal{X}_{\ell-1}^{(j)},{y}^{(j)})\}_{j=1}^{m}$ as training samples, construct a kernel matrix $\mathcal{K}_{\ell}^{k}\in\mathbb{R}^{nm\times nm}$ as in Section 4.2.
3:     Compute a factorization or an approximation of $\mathcal{K}_{\ell}^{k}$ to obtain $Q_{\ell}^{k}\in\mathbb{R}^{nm\times P}$.
4:     For any sample $\mathcal{X}_{\ell-1}^{(j)}$, form a graph signal $Z_{\ell}^{k}(\mathcal{X}_{\ell-1}^{(j)})\in\mathbb{R}^{n\times P}$ via line 6 of Algorithm 1.
5:     Solve (10) via (11) to obtain a the filters $\hat{\mathcal{A}}_{\ell}$.
6:     Compute the output $\mathcal{X}_{\ell}^{(j)}$ of layer $\ell$ via Algorithm 1.
Output: Predictors $\{\hat{\Psi}^{\hat{\mathcal{A}}_{\ell}}(\cdot)\}_{\ell=1}^{L}$ and the output $\mathcal{X}_{L}^{(j)}$.
Algorithm 2 summarizes the training process of a multi-layer CGCN, estimating the parameters of each layer from bottom to top (i.e., in a layer-wise fashion). As such, to learn the nonlinear filter $\tau_{\ell}^{f}$ at layer $\ell$, we should learn the $P$-dimensional vector $\hat{\mathbf{a}}_{\ell k}^{f}$. For this sake, we define the graph signal $Z_{j}^{k}\in\mathbb{R}^{n\times P}$ for any $k\in[K]\cup\{0\}$ and $j\in[m]$ whose $i^{\text{th}}$ row is $[Q^{k}]_{(i,j)}$. We can then apply the nuclear norm relaxation from Section 4.1. Solving the resulting optimization problem yields a matrix of filters $\hat{\mathcal{A}}_{\ell k}\in\mathbb{R}^{P\times F_{\ell}}$ for layer $\ell$ whose $f^{\text{th}}$ column is $\hat{\mathbf{a}}_{\ell k}^{f}$.
Given regularization parameters {}$\{\mathcal{B}_{\ell}\}_{\ell=1}^{L}$ for the filters’ nuclear norms, we thus aim to solve the following optimization problem at each layer $\ell$ due to line 9 of Algorithm 1:
$\hat{\mathcal{A}}_{\ell}\in\operatorname*{arg\,min}_{\|\hat{\mathcal{A}}_{\ell}\|_{*}\leq\mathcal{B}_{\ell}}\tilde{J}(\hat{\mathcal{A}}_{\ell}),\quad\text{ where }\tilde{J}(\hat{\mathcal{A}}_{\ell}):=\frac{1}{m}\sum_{j=1}^{m}J(\hat{\Psi}^{\hat{\mathcal{A}}_{\ell}}(\mathcal{X}_{\ell-1}^{(j)}),y^{(j)}).$ (10)
This can be easily solved via projected gradient descent: at round $t$, with step size $\eta^{t}>0$, compute:
$\hat{\mathcal{A}}_{\ell}^{t+1}=\Pi_{\mathcal{B}_{\ell}}(\hat{\mathcal{A}}_{\ell}^{t}-\eta^{t}\nabla_{\hat{\mathcal{A}}_{\ell}}\tilde{J}(\hat{\mathcal{A}}_{\ell}^{t}))$ (11)
where $\nabla_{\hat{\mathcal{A}}_{\ell}}\tilde{J}$ is the gradient of $\tilde{J}$ from (10) and $\Pi_{\mathcal{B}_{\ell}}$ denotes the Euclidean projection onto the nuclear norm ball $\{\hat{\mathcal{A}}_{\ell}:\|\hat{\mathcal{A}}_{\ell}\|_{*}\leq\mathcal{B}_{\ell}\}$. This projection can be done by first computing the singular value decomposition of $\hat{\mathcal{A}}$ and then projecting the vector of singular values onto the $\ell_{1}$-ball via an efficient algorithm (see, e.g., [13, 12, 50]). The problem in (10) can also be solved using a projected version of any other optimizer (e.g., Adam [23]).
4.4 Scalability, Time Complexity and Challenges of Learning CGCNs
Algorithm 2’s runtime highly depends on how the factorization matrix $Q_{\ell}^{k}$ is computed and its width $P$ (line 3). Naively picking $P=nm$ requires solving the exact kernelized problem, which can be costly. To gain scalability, we compute a Nyström approximation [48] in $\mathcal{O}(P^{2}nm)$ time by randomly sampling training examples, computing their kernel matrix and representing each training example by its similarity to the sampled examples and kernel matrix. This avoids computing or storing the full kernel matrix, thus requiring only $\mathcal{O}(Pmn)$ space. While this approach is efficient [48, 9], it relies on the sampled subset adequately representing the entire data; poorly represented examples may degrade approximation quality. See Appendix E for more details and alternative methods.
4.5 Convergence and Generalization Error
We now analyze Algorithm 2’s generalization error, focusing on models with a single hidden layer ($L=1$) in the binary classification case where the output dimension is $G=1$. Our main result can be easily extended to the multi-class case by a standard one-versus-all reduction to the binary case. In case of multiple hidden layers, our analysis then applies to each hidden layer separately. Specifically, by an approach similar to Zhang et al. [59], we prove that the generalization error of Algorithm 2 converges to the optimal generalization error of a GCN, as it is bounded above by this optimal error plus an additive term that decays polynomially to zero with the sample size. For any $\ell\in[L]$ and $k\in[K]$, we use the random kernel matrix $\mathcal{K}_{\ell}^{k}(\mathcal{X})\in\mathbb{R}^{n\times n}$ induced by a graph signal $\mathcal{X}\in\mathbb{R}^{n\times F}$ drawn randomly from the population, i.e., the $(i,i^{\prime})$-th entry of $\mathcal{K}_{\ell}^{k}(\mathcal{X})$ is $\kappa_{\ell}(\mathbf{z}_{i}^{k}(\mathcal{X}),\mathbf{z}_{i^{\prime}}^{k}(\mathcal{X}))$, where $\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}):=[\mathcal{S}^{k}\mathcal{X}_{\ell-1}]_{i}$ aggregates information at node $i$ from $k$-hops away.
Theorem 4.7.
Consider models with a single hidden layer for binary classification (i.e., $L=G=1$). Let $J(\cdot,\mathbf{y})$ be $M$-Lipchitz continuous for any fixed $\mathbf{y}\in[G]^{n}$ and let $\kappa_{L}$ be either the inverse polynomial kernel or the Gaussian RBF kernel (Recall Appendices B.1-B.2). For any valid activation function $\sigma$ (Recall Appendix B.3), there is a constant radius $R_{L}>0$ such that the expected generalization error is at most:
$\mathbb{E}_{\mathcal{X},\mathbf{y}}[J(\hat{\Psi}^{\hat{\mathcal{A}}_{L}}(\mathcal{X}),\mathbf{y})]\leq\inf_{\Phi\in\mathcal{F}_{\text{gcn}}}\mathbb{E}_{\mathcal{X},\mathbf{y}}[J(\Phi(\mathcal{X}),\mathbf{y})]+\frac{Const}{\sqrt{m}},$
where $Const$ is some constant that depends on $R_{L}$, but not on the sample size $m$.
Proof.
(Sketch) The full proof in Appendix F comprises of two steps: When $L=G=1$, we first consider a wealthier function class, denoted as $\mathcal{F}_{\text{cgcn}}$, which is proven to contain the class of GCNs $\mathcal{F}_{\text{gcn}}$ from (2). The richer class relaxes the class of GCNs in two aspects: (1) The graph filters are relaxed to belong to the RKHS induced by the kernel, and (2) the $\ell_{2}$-norm constraints in (2) are substituted by a single constraint on the Hilbert norm induced by the kernel. Though Algorithm 2 never explicitly optimizes over the function class $\mathcal{F}_{\text{cgcn}}$, we prove that the predictor $\hat{\Psi}^{\hat{\mathcal{A}}_{L}}$ generated at the last layer is an empirical risk minimizer in $\mathcal{F}_{\text{cgcn}}$. Then, we prove that $\mathcal{F}_{\text{cgcn}}$ is not "too big" by upper bounding its Rademacher complexity, a key concept in empirical process theory that can be used to bound the generalization error in our ERM problem (see, e.g., [2] for a brief overview on Rademacher complexity). Combining this upper bound with standard Rademacher complexity theory [2], we infer that the generalization error of $\hat{\Psi}^{\hat{\mathcal{A}}_{L}}$ converges to the best generalization error of $\mathcal{F}_{\text{cgcn}}$. Since $\mathcal{F}_{\text{gcn}}\subset\mathcal{F}_{\text{cgcn}}$, the latter error is bounded by that of GCNs, completing the proof. ∎
5 Experimental Evaluation
Datasets. We evaluate our framework through extensive experiments on several benchmark graph classification datasets [31], including the four bioinformatics datasets MUTAG, PROTEINS, PTC-MR, NCI1 and the chemical compounds dataset Mutagenicity. Following many prior works [29, 53], we randomly split each dataset into three parts: 80% as training set, 10% as validation set and the remaining 10% as test set. The statistics of these datasets appear in Appendix G.1.
Baselines. We apply our convexification procedure to the message-passing mechanisms of popular GNNs and hybrid graph transformers, creating convex counterpartsc. Our framework is integrated into a wide range of leading models, including GCN [7, 24], GAT [45], GATv2 [4], GIN [51], GraphSAGE [20], ResGatedGCN [3], as well as the hybrid transformers GraphGPS [36] and GraphTrans [40, 49]. All models share the same architecture design for fair comparison; see Appendix G.2 for details.
Optimization. All models are trained over $200$ epochs with the Adam optimizer [23] using the cross entropy loss, a starting learning rate of $10^{-3}$, a minimum learning rate of $10^{-6}$ and a weight decay of $10^{-5}$. We use a ReduceLROnPlateau scheduler, which cuts the learning rate in half with a patience of $20$ epochs, and training ends when we reach the minimum learning rate.
Results and Analysis. Table 1 presents our main empirical results, comparing two-layer convex models against two-layer and six-layer non-convex models. Impressively, in most cases, the two-layer convex variants substantially surpass their non-convex counterparts by 10–40% accuracy, showcasing the practical strength of our framework alongside its theoretical benefits. Even when improvements are modest, our convex models match or slightly surpass their non-convex versions, underscoring the robustness and broad applicability of our approach. Notable examples are GATv2 and GraphGPS, whose convex versions significantly outperform their non-convex ones in 4 out of 5 datasets, with slight improvements on Mutagenicity. Our convex models’ superiority over the hybrid transformer GraphGPS, which combines message-passing with self-attention, highlights the power of convexified GNNs over non-convex ones, even when enhanced by self-attention. To visualize these trends more holistically, we also include additional plots in Appendix H.
Over-parameterization is known to aid gradient-based methods to reach global minima that generalize well in non-convex models [1, 11, 60], making it a common practice for obtaining strong performance. Our results challenge this approach. We show that our shallow convex models often outperform their more over-parameterized non-convex counterparts, obtaining better results with fewer layers and parameters. In effect, we beat them twice: in performance and in resource efficiency. These gains are especially notable in small-data regimes, which are prone to overfitting, where we suspect that local optima obtained by non-convex models may hinder generalization. In contrast, convex models avoid such pitfalls, offering more stable and reliable learning. Our findings show that rich representations can be attained without over-parameterization by exploiting convexity in shallow models.
Table 1: Accuracy (%±$\%\pm\sigma$) of two-layer convex versions of base models against their two-layer and six-layer non-convex counterparts on various graph classification datasets over $4$ runs, each with different seeds. The datasets are ordered from left to right in ascending order of size. Convex variants start with ’C’. A two-layer model is marked by (2L), while a six-layer one is marked by (6L). Results of convex models surpassing their non-convex variants by a large gap ($\sim$10-40%) are in dark green and modest improvements in light green.
Model MUTAG PTC-MR PROTEINS NCI1 Mutagenicity
CGCN (2L) $\mathbf{71.2\pm 2.1}$ $\mathbf{59.7\pm 4.2}$ $\mathbf{70.9\pm 2.4}$ $\mathbf{62.3\pm 1.1}$ $\mathbf{61.9\pm 2.7}$
GCN (2L) $55.0\pm 22.0$ $48.5\pm 10.4$ $59.5\pm 4.1$ $50.8\pm 1.7$ $54.7\pm 1.1$
GCN (6L) $61.2\pm 21.6$ $55.0\pm 15.8$ $58.7\pm 6.1$ $50.8\pm 1.7$ $54.7\pm 1.9$
CGIN (2L) $\mathbf{83.2\pm 7.3}$ $\mathbf{63.2\pm 4.7}$ $\mathbf{70.1\pm 2.6}$ $\mathbf{63.6\pm 2.3}$ $\mathbf{59.1\pm 1.8}$
GIN (2L) $60.0\pm 17.6$ $45.0\pm 9.5$ $50.4\pm 12.0$ $55.5\pm 1.4$ $54.8\pm 2.0$
GIN (6L) $61.2\pm 1.6$ $48.5\pm 13.4$ $61.6\pm 1.2$ $60.8\pm 2.8$ $55.8\pm 2.0$
CGAT (2L) $\mathbf{72.5\pm 4.7}$ $\mathbf{59.7\pm 6.0}$ $\mathbf{71.8\pm 2.5}$ $\mathbf{64.6\pm 0.0}$ $\mathbf{58.8\pm 3.6}$
GAT (2L) $65.0\pm 9.3$ $51.4\pm 15.7$ $60.0\pm 3.3$ $51.9\pm 2.2$ $54.8\pm 2.0$
GAT (6L) $68.7\pm 4.1$ $54.2\pm 17.2$ $63.1\pm 1.8$ $52.3\pm 2.9$ $54.8\pm 2.0$
CGATv2 (2L) $\mathbf{68.7\pm 4.1}$ $\mathbf{59.0\pm 5.2}$ $\mathbf{71.3\pm 0.6}$ $\mathbf{63.6\pm 1.1}$ $\mathbf{62.3\pm 4.2}$
GATv2 (2L) $55.0\pm 19.0$ $49.2\pm 15.3$ $61.6\pm 1.6$ $49.5\pm 2.0$ $55.9\pm 1.4$
GATv2 (6L) $57.5\pm 16.0$ $49.2\pm 15.3$ $63.3\pm 2.1$ $49.5\pm 2.0$ $59.6\pm 6.2$
CResGatedGCN (2L) $\mathbf{68.7\pm 4.1}$ $\mathbf{60.4\pm 5.8}$ $\mathbf{72.5\pm 1.0}$ $\mathbf{62.4\pm 3.2}$ $\mathbf{59.6\pm 3.7}$
ResGatedGCN (2L) $57.0\pm 16.7$ $50.7\pm 17.7$ $47.9\pm 9.9$ $56.8\pm 6.6$ $56.1\pm 2.1$
ResGatedGCN (6L) $65.0\pm 15.8$ $50.7\pm 7.1$ $55.8\pm 8.7$ $61.1\pm 3.1$ $56.1\pm 2.1$
CGraphSAGE (2L) $\mathbf{68.7\pm 4.1}$ $\mathbf{65.3\pm 11.5}$ $\mathbf{73.8\pm 3.7}$ $\mathbf{63.7\pm 1.0}$ $\mathbf{66.3\pm 5.3}$
GraphSAGE (2L) $57.4\pm 16.0$ $52.8\pm 7.4$ $62.5\pm 0.8$ $53.1\pm 6.0$ $58.1\pm 3.6$
GraphSAGE (6L) $51.2\pm 1.9$ $55.7\pm 12.5$ $63.1\pm 1.8$ $53.2\pm 6.1$ $60.0\pm 3.3$
CGraphTrans (2L) $\mathbf{68.7\pm 4.1}$ $\mathbf{63.3\pm 10.8}$ $\mathbf{70.7\pm 3.1}$ $\mathbf{63.1\pm 2.8}$ $\mathbf{63.2\pm 4.7}$
GraphTrans (2L) $57.4\pm 16.0$ $40.4\pm 8.0$ $60.7\pm 2.6$ $55.0\pm 3.7$ $53.6\pm 1.3$
GraphTrans (6L) $67.5\pm 7.4$ $49.2\pm 12.3$ $61.6\pm 1.2$ $61.5\pm 2.5$ $59.9\pm 4.6$
CGraphGPS (2L) $\mathbf{83.7\pm 2.1}$ $\mathbf{65.5\pm 9.3}$ $\mathbf{81.6\pm 7.2}$ $\mathbf{60.1\pm 6.1}$ $\mathbf{59.4\pm 4.1}$
GraphGPS (2L) $50.9\pm 19.6$ $49.2\pm 20.1$ $38.0\pm 1.9$ $50.8\pm 2.6$ $41.4\pm 8.7$
GraphGPS (6L) $52.5\pm 2.1$ $46.4\pm 2.3$ $38.8\pm 1.5$ $49.4\pm 1.9$ $55.1\pm 5.5$
6 Conclusions and Future Work
We presented Convexified Message-Passing Graph Neural Networks (CGNNs), a novel framework that transforms the training of message-passing GNNs to a convex optimization problem. They ensure global optimality and computational efficiency, while having provable generalization and statistical guarantees. Empirically, we demonstrated that our CGNNs framework can effectively capture complex patterns without the need for over-parameterization or extensive data. This insight offers a more efficient and data-scarce alternative to traditional deep learning approaches, with substantial improvements across various benchmark datasets. As discussed in Section 4.4, the training of CGNNs is heavily affected by the factorization of the kernel matrix in terms of time and space complexity. Addressing these challenges is thus a key direction for future work.
Appendix A Upper Bounding the Nuclear Norm
Lemma 4.1.
For any layer $\ell$, $\|\mathcal{A}_{\ell}\|_{*}\leq\mathcal{B}_{\ell}$ for some constant $\mathcal{B}_{\ell}>0$ dependent on $R_{\ell},F_{\ell},F_{\ell-1},K$.
Proof.
Note that the nuclear norm of any matrix $B$ can be bounded as $\|B\|_{*}\leq\|B\|_{F}\sqrt{\operatorname*{rank}(B)}$, where $\|B\|_{F}$ is the Frobenius norm which is identical to the standard Euclidean norm of the vectorized version of the matrix. Hence, the nuclear norm of the filters $\mathcal{A}_{\ell}$ satisfying (5a) and (5b) is upper bounded by
$\|\mathcal{A}_{\ell}\|_{*}^{2}\leq\|\mathcal{A}_{\ell}\|_{F}^{2}F_{\ell-1}={\sum_{k=0}^{K}\sum_{f\in[F_{\ell}]}\sum_{g\in[F_{\ell-1}]}|a_{\ell k}^{fg}|^{2}}\cdot{F_{\ell-1}}\leq R_{\ell}^{2}{F_{\ell}F_{\ell-1}(K+1)}.$
That is, $\|\mathcal{A}_{\ell}\|_{*}\leq\mathcal{B}_{\ell}$ for $\mathcal{B}_{\ell}:=R_{\ell}\sqrt{F_{\ell}F_{\ell-1}(K+1)}$. ∎
Appendix B Choice of Kernels and Valid Activations
In this section, we provide properties regarding two types of kernels: the inverse polynomial kernel (Section B.1) and the Gaussian RBF kernel (Section B.2). We show that the reproducing kernel Hilbert space (RKHS) corresponding to these kernels contain the nonlinear filter at each layer $\ell$ for the $f^{\text{th}}$ feature of node $i$ for certain choices of an activation function $\sigma$. Finally, we summarize the possible choices for an activation function $\sigma$ that result from our discussion (Section B.3).
B.1 Inverse Polynomial Kernel
For $F$ features, the inverse polynomial kernel $\kappa^{IP}:\mathbb{R}^{F}\times\mathbb{R}^{F}\rightarrow\mathbb{R}$ is given as follows:
$\kappa^{IP}(\mathbf{z},\mathbf{z}^{\prime}):=\frac{1}{2-\langle\mathbf{z},\mathbf{z}^{\prime}\rangle};\quad\|\mathbf{z}\|_{2}\leq 1,\|\mathbf{z}^{\prime}\|_{2}\leq 1$ (12)
Zhang et al. [59] prove that $\kappa^{IP}$ indeed constitutes a legal kernel function. We include the proof to make the paper self-contained: they provide a feature mapping $\varphi:\mathbb{R}^{F}\rightarrow\ell^{2}(\mathbb{N})$ for which $\kappa^{IP}(\mathbf{z},\mathbf{z}^{\prime})=\langle\varphi(\mathbf{z}),\varphi(\mathbf{z}^{\prime})\rangle$. The $(k_{1},\dots,k_{j})$-th coordinate of $\varphi(\mathbf{z})$, where $j\in\mathbb{N}$ and $k_{1},\dots,k_{j}\in[F_{\ell}]$, is given by $2^{-\frac{j+1}{2}}x_{k_{1}}\dots x_{k_{j}}$. Thus:
$\langle\varphi(\mathbf{z}),\varphi(\mathbf{z}^{\prime})\rangle=\sum_{j=0}^{\infty}2^{-(j+1)}\sum_{(k_{1},\dots,k_{j})\in[F_{\ell}]^{j}}z_{k_{1}}\dots z_{k_{j}}z^{\prime}_{k_{1}}\dots z^{\prime}_{k_{j}}$ (13)
Since $\|\mathbf{z}\|_{2}\leq 1$ and $\|\mathbf{z}^{\prime}\|_{2}\leq 1$, the series above is absolutely convergent. Combined with the fact that $|\langle\mathbf{z},\mathbf{z}^{\prime}\rangle|\leq 1$, this yields a simplified form of (13):
$\langle\varphi(\mathbf{z}),\varphi(\mathbf{z}^{\prime})\rangle=\sum_{j=0}^{\infty}2^{-(j+1)}(\langle\mathbf{z},\mathbf{z}^{\prime}\rangle)^{j}=\frac{1}{2-\langle\mathbf{z},\mathbf{z}^{\prime}\rangle}=\kappa^{IP}(\mathbf{z},\mathbf{z}^{\prime})$ (14)
as desired. Combining (7) with the proof of Lemma 1 in Appendix A.1 of Zhang et al. [58], the RKHS associated with $\kappa^{IP}$ is comprised of the class of nonlinear graph filters in (7). It can be formulated as follows:
Corollary B.1.
Assume that $\sigma$ has a polynomial expansion $\sigma(t)=\sum_{j=0}^{\infty}a_{j}t^{j}$. Let $C_{\sigma}(t):=\sqrt{\sum_{j=0}^{\infty}2^{j+1}a_{j}^{2}t^{2j}}$. If $C_{\sigma}(\|\mathbf{a}^{f}_{\ell k}\|_{2})<\infty$, then the RKHS induced by $\kappa^{IP}$ contains the function $\tau_{\ell}^{f}:\mathcal{Z}\in\mathbb{R}^{(K+1)\times F_{\ell-1}}\mapsto\sigma(\sum_{k=0}^{K}\langle[\mathcal{Z}]_{k+1},\mathbf{a}^{f}_{\ell k}\rangle)$ with Hilbert norm $\|\tau_{\ell}^{f}\|_{\mathcal{A}}\leq C_{\sigma}(\sum_{k=0}^{K}\|\mathbf{a}^{f}_{\ell k}\|_{2})$.
Consequently, Zhang et al. [58] state that upper bounding $C_{\sigma}(t)$ for a particular activation function $\sigma$ is sufficient. To enforce $C_{\sigma}(t)<\infty$, the scalars $\{a_{j}\}_{j=0}^{\infty}$ must rapidly converge to zero, implying that $\sigma$ must be sufficiently smooth. For polynomial functions of degree $d$, the definition of $C_{\sigma}$ yields $C_{\sigma}(t)=\mathcal{O}(t^{d})$. For the sinusoid activation $\sigma(t):=sin(t)$, we have:
$C_{\sigma}(t)=\sqrt{\sum_{j=0}^{\infty}\frac{2^{2j+2}}{((2j+1)!)^{2}}(t^{2})^{2j+1}}\leq 2e^{t^{2}}$ (15)
[…]
Remark B.2.
Generally, a function $f$ is sufficiently smooth if it is differentiable sufficient number of times. For instance, if we were to consider $\sigma(t):=\sin(t)$, its taylor expansion is provided by $\sin(t)=\sum_{j=0}^{\infty}\frac{(-1)^{j}}{(2j+1)!}t^{2j+1}$. Denoting $b_{j}:=\frac{(-1)^{j}}{(2j+1)!}t^{2j+1}$, the ratio test yields $\lim_{j\rightarrow\infty}|\frac{b_{j+1}}{b_{j}}|=0$ for all $t\in I\!R$. Hence, the radius of convergence of the expansion is the set of all real numbers.
Regarding Corollary B.1, as well as Corollary B.3 which follows in the next subsection, for enforcing $C_{\sigma}(t)<\infty$, the coefficients $\{a_{j}\}_{j=0}^{\infty}$ in $\sigma$’s polynomial expansion must quickly converge to zero as illustrated above, thus requiring $\sigma$ to be sufficiently smooth. Such activations include polynomial functions, the sinusoid, the erf function and the smoothed hinge loss.
B.2 Gaussian RBF Kernel
For $F$ features, the Gaussian RBF kernel $\kappa^{RBF}:\mathbb{R}^{F}\times\mathbb{R}^{F}\rightarrow\mathbb{R}$ is given as follows:
$\kappa^{RBF}(\mathbf{z},\mathbf{z}^{\prime}):=e^{-\gamma\|\mathbf{z}-\mathbf{z}^{\prime}\|_{2}^{2}};\quad\|\mathbf{z}\|_{2}=\|\mathbf{z}^{\prime}\|_{2}=1$ (16)
Combining (7) with the proof of Lemma 2 in Appendix A.2 of Zhang et al. [58], the RKHS associated with $\kappa^{RBF}$ is comprised of the class of nonlinear input features. It can be formulated as follows:
Corollary B.3.
Assume that $\sigma$ has a polynomial expansion $\sigma(t)=\sum_{j=0}^{\infty}a_{j}t^{j}$. Let $C_{\sigma}(t):=\sqrt{\sum_{j=0}^{\infty}\frac{j!e^{2\gamma}}{(2\gamma)^{j}}a_{j}^{2}t^{2j}}$. If $C_{\sigma}(\|\mathbf{a}^{f}_{\ell k}\|_{2})<\infty$, then the RKHS induced by $\kappa^{RBF}$ contains the function $\tau_{\ell}^{f}:\mathcal{Z}\in\mathbb{R}^{(K+1)\times F_{\ell-1}}\mapsto\sigma(\sum_{k=0}^{K}\langle[\mathcal{Z}]_{k+1},\mathbf{a}^{f}_{\ell k}\rangle)$ with Hilbert norm $\|\tau_{\ell}^{f}\|_{\mathcal{A}}\leq C_{\sigma}(\sum_{k=0}^{K}\|\mathbf{a}^{f}_{\ell k}\|_{2})$.
Comparing Corollary B.1 and Corollary B.3, $\kappa^{RBF}$ requires a stronger condition on the smoothness of the activation function. For polynomial functions of degree $d$, $C_{\sigma}(t)=\mathcal{O}(t^{d})$ remains intact. However, for the sinusoid activation $\sigma(t):=\sin(t)$, we have:
$C_{\sigma}(t)=\sqrt{e^{2\gamma}\sum_{j=0}^{\infty}\frac{1}{(2j+1)!}(\frac{t^{2}}{2\gamma})^{2j+1}}\leq e^{\frac{t^{2}}{4\gamma}+\gamma}$ (17)
[…]
B.3 Valid Activation Functions
Now, we summarize the possible valid choices of an activation function $\sigma$:
  1. 1.
    Arbitrary polynomial functions.
  2. 2.
    The sinusoid activation function $\sigma(t):=\sin(t)$.
  3. 3.
    The erf function $\sigma_{erf}(t):=\frac{2}{\sqrt{\pi}}\int_{0}^{t}e^{\tau^{2}}\mathrm{d}\tau$, which approximates the sigmoid function.
  4. 4.
    The smoothed hinge loss function $\sigma_{sh}(t):=\int_{-\infty}^{t}\frac{1}{2}(\sigma_{erf}(\tau)+1)\mathrm{d}\tau$, which approximates the ReLU function.
Recall that the erf and smoothed hinge loss activation functions are unsuited for the RBF kernel.
Appendix C Rephrasing the Nonlinear Filter at each Layer
Proof.
For any graph signal $\mathcal{S}^{k}\mathcal{X}$, let $\mathbf{z}_{i}^{k}(\mathcal{X}):=[\mathcal{S}^{k}\mathcal{X}]_{i}$ be the $i^{\text{th}}$ row of $\mathcal{S}^{k}\mathcal{X}$, which is the information aggregated at node $i$ from $k$-hops away. We also denote the $f^{\text{th}}$ column of the filters $\mathcal{A}_{\ell k}$ as $\mathbf{a}^{f}_{\ell k}=[a_{\ell k}^{fg}]_{g\in[F_{\ell-1}]}$. By (1), the value of the $f^{\text{th}}$ feature of node $i$ at layer $\ell$ can be thus written as:
=${}_{if}=\sum_{k=0}^{K}\sum_{g=1}^{F_{\ell-1}}[\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1})]_{g}[\mathbf{a}^{f}_{\ell k}]_{g}=\sum_{k=0}^{K}\langle\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}),\mathbf{a}^{f}_{\ell k}\rangle.$ (18)
Defining the matrix $\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1})\in\mathbb{R}^{(K+1)\times F_{\ell-1}}$ whose $(k+1)^{\text{th}}$ row is $[\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1})]_{k+1}=\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1})$ for $k\in[K]\cup\{0\}$, ∎
Appendix D Kernelization for Nonlinear Activations
In this section, we elaborate on the kernelization for the case of nonlinear activation functions, briefly discussed in Section 4.2. Recall that $\sigma$ denotes, for brevity, its entrywise application in (2). Further, the of the $f^{\text{th}}$ feature of node $i$ at layer $\ell$’s output is given by $\sigma(\sum_{k=0}^{K}\langle\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}),\mathbf{a}^{f}_{\ell k}\rangle)$ by (18). According to Appendix B, consider a sufficiently smooth valid activation function $\sigma$, and a proper choice of kernel $\kappa_{\ell}:\mathbb{R}^{F_{\ell-1}}\times\mathbb{R}^{F_{\ell-1}}\rightarrow\mathbb{R}$. Without loss of generality, we assume that, each input input graph signal $\mathcal{X}^{(j)}$, the feature vector $[\mathcal{X}^{(j)}]_{i}=\mathbf{x}_{i}^{(j)}$ of each node $i$ resides in the unit $\ell_{2}$-ball (which can be obtain via normalization). Combined with $\kappa_{\ell}$’s being a continuous kernel with $\kappa_{\ell}(\mathbf{z},\mathbf{z})\leq 1$, Mercer’s theorem [44, Theorem 4.49] implies that there exists a feature mapping $\varphi_{\ell}:\mathbb{R}^{F_{\ell-1}}\rightarrow\ell^{2}(\mathbb{N})$ for which $\kappa_{\ell}(\mathbf{z},\mathbf{z}^{\prime})=\langle\varphi_{\ell}(\mathbf{z}),\varphi_{\ell}(\mathbf{z}^{\prime})\rangle$, which yields the following for each feature $f$ of any node $i$ at layer $\ell$:
$\sigma\Bigg{(}\sum_{k=0}^{K}\langle\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}),\mathbf{a}^{f}_{\ell k}\rangle\Bigg{)}=\sum_{k=0}^{K}\langle\varphi_{\ell}(\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1})),\varphi_{\ell}(\mathbf{a}^{f}_{\ell k})\rangle=:\sum_{k=0}^{K}\langle\varphi_{\ell}(\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1})),\bar{\mathbf{a}}^{f}_{\ell k}\rangle$ (19)
where $\bar{\mathbf{a}}^{f}_{\ell k}\in\ell^{2}(\mathbb{N})$ is a countable-dimensional vector, due to the fact that $\varphi$ itself is a countable sequence of functions. By Corollary B.1 and Corollary B.3, we have that $\|\bar{\mathbf{a}}^{f}_{\ell k}\|_{2}\leq C_{\sigma,\ell}(\|{\mathbf{a}}^{f}_{\ell k}\|_{2})$, provided a monotonically increasing $C_{\sigma}$ which depends on the kernel $\kappa_{\ell}$. Accordingly, $\varphi_{\ell}(\mathbf{z})$ may be utilized as the vectorized representation of each output features $\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1})$, where $\bar{\mathbf{a}}^{f}_{\ell k}$ constitutes the linear graph filters, thus reducing the problem to training a GCN with the identity activation function.
Each graph filter is now parametrized by the countable-dimensional vector $\bar{\mathbf{a}}^{f}_{\ell k}$. Subsequently, we supply a reduction of the original ERM problem to a finite-dimensional one. Output on the training data $\mathcal{T}:=\{(\mathcal{G}^{(j)},\mathcal{X}^{(j)},y^{(j)})\}_{j=1}^{m}$ should be solely considered when solving the ERM problem, i.e., $\sum_{k=0}^{K}\langle\varphi_{\ell}(\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1})),\bar{\mathbf{a}}^{f}_{\ell k}\rangle$ for each feature $f$ of any node $i$ at layer $\ell$. Let $\mathcal{P}_{\ell-1}^{k}$ be the orthogonal projector onto the linear subspace spanned by $\Gamma_{\ell-1}^{k}:=\{\mathbf{z}_{i}^{k}(\mathcal{X}^{(j)}_{\ell-1}):i\in[n],j\in[m],k\in[K]\cup\{0\}\}$. Therefore, for any $(i,j)\in[n]\times[m]$:
$\sum_{k=0}^{K}\langle\varphi_{\ell}(\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}^{(j)})),\bar{\mathbf{a}}^{f}_{\ell k}\rangle=\sum_{k=0}^{K}\langle\mathcal{P}_{\ell-1}^{k}\varphi_{\ell}(\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}^{(j)})),\bar{\mathbf{a}}^{f}_{\ell k}\rangle=\sum_{k=0}^{K}\langle\varphi_{\ell}(\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}^{(j)})),\mathcal{P}_{\ell-1}^{k}\bar{\mathbf{a}}^{f}_{\ell k}\rangle$ (20)
where the last step stems from the fact that the orthogonal projector $\mathcal{P}_{\ell-1}^{k}$ is self-adjoint. Without loss of generality, for the sake of solving the ERM, we assume that $\bar{\mathbf{a}}^{f}_{\ell k}$ belongs to $\Gamma_{\ell-1}^{k}$. Similarly to (8), we can thus reparameterize it by:
$\bar{\mathbf{a}}^{f}_{\ell k}=\sum_{(i,j)\in[n]\times[m]}\beta_{\ell k,(i,j)}^{f}\varphi_{\ell}(\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}^{(j)}))$ (21)
where we denote $\mathbf{\beta}_{\ell k}^{f}\in\mathbb{R}^{nm}$ as a vector of the coefficients, whose $(i,j)$-th coordinate is $\beta_{\ell k,(i,j)}^{f}$. To estimate $\bar{\mathbf{a}}^{f}_{\ell k}$, it thus suffices to estimate the vector $\mathbf{\beta}_{\ell k}^{f}$. By definition, the relation $(\mathbf{\beta}_{\ell k}^{f})^{T}\mathcal{K}_{\ell}^{k}\mathbf{\beta}_{\ell k}^{f}=\|\bar{\mathbf{a}}^{f}_{\ell k}\|_{2}^{2}$ holds, where $\mathcal{K}_{\ell}^{k}\in\mathbb{R}^{nm\times nm}$ is the symmetric kernel matrix from Section 4.2, whose rows and columns are indexed by some triple $(i,j)\in[n]\times[m]$. The entry at row $(i,j)$ and column $(i^{\prime},j^{\prime})$ equals to $\kappa_{\ell}(\mathbf{z}_{i}^{k}(\mathcal{X}^{(j)}_{\ell-1}),\mathbf{z}_{i^{\prime}}^{k}(\mathcal{X}^{(j^{\prime})}_{\ell-1}))$. A suitable factorization of $\mathcal{K}_{\ell}^{k}$ such that $\mathcal{K}_{\ell}^{k}=Q_{\ell}^{k}(Q_{\ell}^{k})^{\top}$ for $Q_{\ell}^{k}\in\mathbb{R}^{nm\times P}$ then yields the following norm constraint:
$\|(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f}\|_{2}=\sqrt{(\mathbf{\beta}_{\ell k}^{f})^{T}\mathcal{K}_{\ell}^{k}\mathbf{\beta}_{\ell k}^{f}}=\|\bar{\mathbf{a}}^{f}_{\ell k}\|_{2}^{2}\leq C_{\sigma,\ell}(\|{\mathbf{a}}^{f}_{\ell k}\|_{2})\leq C_{\sigma,\ell}(R_{\ell})$ (22)
where the last inequality is due to $\|{\mathbf{a}}^{f}_{\ell k}\|_{2}\leq R_{\ell}$ due to (3).
Let $\mathbf{v}_{\ell}^{k}(\mathbf{z})\in\mathbb{R}^{nm}$ be a vector whose $(i,j)$-th coordinate equals to $\kappa_{\ell}(\mathbf{z},\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}^{(j)}))$. By (19) and (21), we can write:
$\sigma\Bigg{(}\sum_{k=0}^{K}\langle\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}^{(j)}),\mathbf{a}^{f}_{\ell k}\rangle\Bigg{)}\equiv\sum_{k=0}^{K}\langle\varphi_{\ell}(\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}^{(j)})),\bar{\mathbf{a}}^{f}_{\ell k}\rangle\equiv\sum_{k=0}^{K}\langle\mathbf{v}_{\ell}^{k}(\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}^{(j)})),\mathbf{\beta}_{\ell k}^{f}\rangle$ (23)
$\equiv\sum_{k=0}^{K}\langle\mathbf{\beta}_{\ell k}^{f},\mathbf{v}_{\ell}^{k}(\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}^{(j)}))\rangle$
For each $\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}^{(j)})$, the vector $\mathbf{v}_{\ell}^{k}(\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}^{(j)}))$ resides in the column space of the symmetric kernel matrix $\mathcal{K}_{\ell}^{k}$. Thus, given that $(Q_{\ell}^{k})^{\dagger}$ denotes the pseudo-inverse of $Q_{\ell}^{k}$, we infer the following for any $(i,j)\in[n]\times[m]$:
$\sum_{k=0}^{K}\langle\mathbf{\beta}_{\ell k}^{f},\mathbf{v}_{\ell}^{k}(\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}^{(j)}))\rangle=\sum_{k=0}^{K}(\mathbf{\beta}_{\ell k}^{f})^{\top}Q_{\ell}^{k}(Q_{\ell}^{k})^{\dagger}\mathbf{v}_{\ell}^{k}(\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}^{(j)}))$ (24)
$=\sum_{k=0}^{K}\Big{\langle}((Q_{\ell}^{k})^{\top})^{\dagger}(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f},\mathbf{v}_{\ell}^{k}(\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}^{(j)}))\Big{\rangle}$
Accordingly, replacing the vector $\mathbf{\beta}_{\ell k}^{f}$ on the right-hand side of (23) by $((Q_{\ell}^{k})^{\top})^{\dagger}(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f}$ won’t affect the empirical risk. For any matrix $\mathcal{Z}\in\mathbb{R}^{(K+1)\times F_{\ell-1}}$ whose $(k+1)^{\text{th}}$ row is given by $[\mathcal{Z}]_{k+1}:=\mathbf{z}_{k}$ for $k\in[K]\cup\{0\}$, we thereby propose the following parameterization function:
$\tau_{\ell}^{f}(\mathcal{Z}):=\sum_{k=0}^{K}\Big{\langle}((Q_{\ell}^{k})^{\top})^{\dagger}(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f},\mathbf{v}_{\ell}^{k}(\mathbf{z}_{k})\Big{\rangle}=\sum_{k=0}^{K}\langle(Q_{\ell}^{k})^{\dagger}\mathbf{v}_{\ell}^{k}(\mathbf{z}_{k}),(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f}\rangle$ (25)
Consequently, the value of the $f^{\text{th}}$ feature of node $i$ at layer $\ell$ is given by $\tau_{\ell}^{f}(\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1}))$, where $\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1})\in\mathbb{R}^{(K+1)\times F_{\ell-1}}$ is the matrix whose $(k+1)^{\text{th}}$ row is $[\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1})]_{k}=\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1})$ for $k\in[K]\cup\{0\}$. That is, $(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f}$ now act as the filters under the reparameterization in (25) which satisfy the following constraints:
Norm Constraint. $\|(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f}\|_{2}\leq C_{\sigma,\ell}(R_{\ell})$ due to (22).
Rank Constraint. Its rank is at most $F_{\ell-1}$.
Therefore, similarly to the linear case in Section 4.1, those constraints can be relaxed to the nuclear norm constraint:
$\|(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f}\|_{*}\leq C_{\sigma,\ell}(R_{\ell})\sqrt{F_{\ell}F_{\ell-1}(K+1)}$ (26)
Comparing the constraint (6) from the linear activation function case with (26), we notice that the sole distinction is that the term $R_{\ell}$ is replaced by $C_{\sigma,\ell}(R_{\ell})$, which is required due to our use of the kernel trick for reducing the case of nonlinear activation functions to the linear setting.
Appendix E More Details on Matrix Approximation and Factorization Methods
As mentioned in Section 4.4, the time complexity of Algorithm 2 largely depends on the width $P$ of each factorization matrix $Q_{\ell}^{k}$ and the method used to obtain it (line 3). A naive choice of $P=nm$ corresponds to solving the full kernelized problem, which can be computationally expensive. For instance, using Cholesky factorization with $P=nm$ [8] incurs a complexity of $\mathcal{O}(m^{3}n^{3})$ and space complexity of $\mathcal{O}(m^{2}n^{2})$.
However, to gain scalability, we use Nyström approximation [48]: an approximation $\mathcal{K}_{\ell}^{k}\approx Q_{\ell}^{k}(Q_{\ell}^{k})^{\top}$ for $Q_{\ell}^{k}\in\mathbb{R}^{nm\times P}$ is obtained by randomly sampling $P$ rows/columns from the original $\mathcal{K}_{\ell}^{k}$, which takes $\mathcal{O}(P^{2}nm)$ time. In practice, it randomly samples training examples, computes their kernel matrix and represents each training example by its similarity to the sampled examples and kernel matrix. It is well-known that this method can significantly accelerate our computations [48, 9], which is attained by using the approximate kernel matrix instead of the full one. Another advantage is that it is not necessary to compute or store the full kernel matrix, but only a submatrix requiring space complexity of $\mathcal{O}(Pmn)$.
The Nyström method works effectively so long as the sampled training examples and the kernel matrix adequately represent the entire graph. However, if some training examples are distant from the randomly sampled ones, they may not be well-represented.
Another possible approximation method is the random feature approximation [35], which can be executed in $\mathcal{O}(nmP\log F_{\ell})$ time [26]. Overall, using approximations significantly accelerates our algorithm’s performance in terms of time and space complexity.
Appendix F Proof of Theorem 4.7
Theorem 4.7.
Consider models with a single hidden layer for binary classification (i.e., $L=G=1$). Let $J(\cdot,\mathbf{y})$ be $M$-Lipchitz continuous for any fixed $\mathbf{y}\in[G]^{n}$ and let $\kappa_{L}$ be either the inverse polynomial kernel or the Gaussian RBF kernel (Recall Appendices B.1-B.2). For any valid activation function $\sigma$ (Recall Appendix B.3), there is a constant $C_{\sigma,L}(\mathcal{B}_{L})$ such that with the radius $R_{L}:=C_{\sigma,L}(\mathcal{B}_{L})\sqrt{F}$, the expected generalization error is at most:
$\mathbb{E}_{\mathcal{X},\mathbf{y}}[J(\hat{\Psi}^{\hat{\mathcal{A}}_{L}}(\mathcal{X}),\mathbf{y})]\leq\inf_{\Phi\in\mathcal{F}_{\text{gcn}}}\mathbb{E}_{\mathcal{X},\mathbf{y}}[J(\Phi(\mathcal{X}),\mathbf{y})]+\frac{cMR_{L}\sqrt{\log(n(K+1))\sum_{k=0}^{K}\mathbb{E}[||K_{L}^{k}(\mathcal{X})||_{2}]}}{\sqrt{m}}$ (27)
where $c>0$ is a universal constant.
We begin by proving several properties that apply to the general multi-class case with models comprising of multiple hidden layers, and then we specialize them to a single-hidden-layer models in the binary classification case. First, a wealthier function class shall be considered, which contains the class of GCNs. For each layer $\ell$, let $\mathcal{H}_{\ell}$ be the RKHS induced by the kernel function $\kappa_{\ell}$ and $||\cdot||_{\mathcal{H}_{\ell}}$ be the associated Hilbert norm. Recall that, given any matrix $\mathcal{Z}\in\mathbb{R}^{(K+1)\times F_{\ell-1}}$, we denote its $(k+1)^{\text{th}}$ row as $[\mathcal{Z}]_{k+1}:=\mathbf{z}_{k}$ for any $k\in[K]\cup\{0\}$. Using this notation, consider the following function class:
$\mathcal{F}_{\text{cgcn}}:=\{\Phi^{\mathcal{A}}=\phi^{{\mathcal{A}}_{L}}\circ\dots\circ\phi^{{\mathcal{A}}_{1}}:\phi^{{\mathcal{A}}_{\ell}}\in\mathcal{F}_{\ell,cgcn}\quad\forall\ell\in[L]\}\text{, where}$ (28)
$\mathcal{F}_{\ell,cgcn}:=\Bigg{\{}\phi^{{\mathcal{A}}_{\ell}}:\mathbb{R}^{n\times F_{\ell-1}^{\star}}\rightarrow\mathbb{R}^{n\times F_{\ell}}:F_{\ell-1}^{\star}<\infty\text{ and }[\phi^{{\mathcal{A}}_{\ell}}(\mathcal{X})]_{if}=\tau_{\ell}^{f}(\mathcal{Z}_{i,\ell}(\mathcal{X}))$
[…]
$\|\tau_{\ell}^{f}\|_{\mathcal{H}_{\ell}}\leq C_{\sigma,\ell}(R_{\ell})(K+1)\sqrt{F_{\ell}F_{\ell-1}}\Bigg{\}}$
where $C_{\sigma,\ell}(R_{\ell})$ depends solely on the chosen activation function $\sigma$ and the regularization parameter $R_{\ell}$. We further consider the function class of all graph filters at layer $\ell$ of a standard GCN:
$\mathcal{F}_{\ell,gcn}:=\Bigg{\{}\Psi^{\mathcal{A}_{\ell}}:\Psi^{\mathcal{A}_{\ell}}(\mathcal{X}):=\sigma\Bigg{(}\sum_{k=0}^{K}\mathcal{S}^{k}\mathcal{X}_{\ell-1}\mathcal{A}_{\ell k}\Bigg{)}\text{ and }\max_{0\leq k\leq K}\|\mathbf{a}^{f}_{\ell k}\|_{2}\leq R_{\ell}\text{ }\forall f\Bigg{\}}$ (29)
Next, we show that the class $\mathcal{F}_{\text{cgcn}}$ contains the class of GCNs $\mathcal{F}_{\text{gcn}}$, making the former a reacher class of functions. Particularly, the class $\mathcal{F}_{\ell,cgcn}$ of kernelized filters at layer $\ell$ contains the class $\mathcal{F}_{\ell,gcn}$ of all graph filters at layer $\ell$ corresponding to a standard GCN.
Lemma F.1.
For any valid activation function $\sigma$, there is some $C_{\sigma,\ell}(R_{\ell})$ that solely depends on $\sigma$ and $R_{\ell}$ such that $\mathcal{F}_{\text{gcn}}\subset\mathcal{F}_{\text{cgcn}}$. In particular, $\mathcal{F}_{\ell,gcn}\subset\mathcal{F}_{\ell,cgcn}$.
Proof.
Recall that that any GCN $\Phi\in\mathcal{F}_{\text{gcn}}$ induces a nonlinear filter at each layer $\ell$ for the $f^{\text{th}}$ which can be characterized by $\tau_{\ell}^{f}:\mathcal{Z}\in\mathbb{R}^{(K+1)\times F_{\ell-1}}\mapsto\sigma(\sum_{k=0}^{K}\langle[\mathcal{Z}]_{k+1},\mathbf{a}^{f}_{\ell k}\rangle)$, i.e., the GCN $\Phi$ is given by the following composition of functions $\Phi=\tau_{L}\circ\dots\circ\tau_{1}$, where we defined $\tau_{\ell}:\mathbf{z}\mapsto(\tau_{\ell}^{f}(\mathbf{z}))_{f\in[F_{\ell}]}$. Given any valid activation function, Corollary B.1 and Corollary B.3 yield that $\tau_{\ell}^{f}$ resides within the reproducing kernel Hilbert space $\mathcal{H}_{\ell}$ and its Hilbert norm is upper bounded as $\|\tau_{\ell}^{f}\|_{\mathcal{H}_{\ell}}\leq C_{\sigma}(\|\sum_{k=0}^{K}\mathbf{a}^{f}_{\ell k}\|_{2})\leq C_{\sigma}(R_{\ell})\sqrt{K+1}$, which clearly satisfies the constraint in (28) for any layer $\ell$, as desired. ∎
Recalling that $\hat{\Psi}^{\hat{\mathcal{A}}_{\ell}}(\cdot)$ is the predictor at each layer $\ell$ produced by Algorithm 2, we now formally prove that the predictor $\hat{\Psi}^{\hat{\mathcal{A}}_{L}}(\cdot)$ generated at the last layer is an empirical risk minimizer in the function class $\mathcal{F}_{\text{cgcn}}$.
Lemma F.2.
By running Algorithm 2 with the regularization parameter $\mathcal{B}_{\ell}=C_{\sigma,\ell}(R_{\ell})\sqrt{F_{\ell}F_{\ell-1}}$ at layer $\ell$, the predictor $\hat{\Psi}^{\hat{\mathcal{A}}_{L}}(\cdot)$ generated at the last layer satisfies:
$\hat{\Psi}^{\hat{\mathcal{A}}_{L}}\in\operatorname*{arg\,min}_{\Phi\in\mathcal{F}_{\text{cgcn}}}\frac{1}{m}\sum_{j=1}^{m}J(\Phi(\mathcal{X}^{(j)}),y^{(j)})$ (30)
Proof.
Consider the function class of all CGCNs whose filters at layer $\ell$ have a nuclear norm of at most $\mathcal{B}_{\ell}$, given by:
:=$\mathcal{F}_{\mathcal{B}_{\ell}}:=\Bigg{\{}\Phi^{\mathcal{A}}=\Psi^{\mathcal{A}_{L}}\circ\dots\circ\Psi^{\mathcal{A}_{1}}\Bigg{|}$ $\Psi^{\mathcal{A}_{\ell}}:\mathcal{X}\mapsto\sum_{k=0}^{K}\mathcal{S}^{k}\mathcal{X}_{\ell-1}\mathcal{A}_{\ell k}\text{ and }$ (31)
$\mathcal{A}_{\ell k}\in\mathbb{R}^{F_{\ell-1}\times F_{\ell}}\text{ and }\mathcal{A}_{\ell}\|_{*}\leq\mathcal{B}_{\ell}\Bigg{\}}$
Our goal is thus proving that $\mathcal{F}_{\mathcal{B}_{\ell}}\subset\mathcal{F}_{\text{cgcn}}$, and that any empirical risk minimizer in $\mathcal{F}_{\text{cgcn}}$ is also in $\mathcal{F}_{\mathcal{B}_{\ell}}$. We begin with showing that $\mathcal{F}_{\mathcal{B}_{\ell}}\subset\mathcal{F}_{\text{cgcn}}$. Let $\Phi^{\mathcal{A}}=\Psi^{\mathcal{A}_{L}}\circ\dots\circ\Psi^{\mathcal{A}_{1}}\in\mathcal{F}_{\mathcal{B}_{\ell}}$ such that $\Psi^{\mathcal{A}_{\ell}}(\mathcal{X})=\sum_{k=0}^{K}\mathcal{S}^{k}\mathcal{X}_{\ell-1}\mathcal{A}_{\ell k}$ with $\|\mathcal{A}_{\ell}\|_{*}\leq\mathcal{B}_{\ell}$ for any $\ell\in[L]$. We shall prove that $\Psi^{\mathcal{A}_{\ell}}\in\mathcal{F}_{\ell,cgcn}$. Indeed, consider the singular value decomposition (SVD) of $\mathcal{A}_{\ell k}$, which we assume is given by $\mathcal{A}_{\ell k}=\sum_{s\in[F_{\ell-1}^{\star}]}\lambda_{\ell ks}\mathbf{w}_{\ell ks}\mathbf{u}_{\ell ks}^{\top}$ for some $F_{\ell-1}^{\star}<\infty$, where each left-singular vector $\mathbf{w}_{\ell ks}$ and each right-singular vector $\mathbf{u}_{\ell ks}$ are both unit column vectors, while each singular value $\lambda_{\ell ks}$ is a real number. Therefore, the filters corresponding to the $f^{\text{th}}$ feature can be written as:
$\mathbf{a}^{f}_{\ell k}=\sum_{s\in[F_{\ell-1}^{\star}]}\lambda_{\ell ks}[\mathbf{w}_{\ell ks}]_{f}\cdot\mathbf{u}_{\ell ks}^{\top}=\sum_{s\in[F_{\ell-1}^{\star}]}\lambda_{\ell ks}w_{\ell ks}^{f}\cdot\mathbf{u}_{\ell ks}^{\top}$ (32)
As such, the value of the $f^{\text{th}}$ feature of node $i$ under $\Psi^{\mathcal{A}_{\ell}}$ can be rephrased as:
=${}_{if}=\sum_{k=0}^{K}\sum_{g=1}^{F_{\ell-1}}[\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1})]_{g}[\mathbf{a}^{f}_{\ell k}]_{g}=\sum_{k=0}^{K}\Bigg{\langle}\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}),\sum_{s\in[F_{\ell-1}^{\star}]}\lambda_{\ell ks}w_{\ell ks}^{f}\mathbf{u}_{\ell ks}^{\top}\Bigg{\rangle}$ (33)
Let $\mathbf{v}_{\ell}^{k}(\mathbf{z})\in\mathbb{R}^{nm}$ be a vector whose $(i,j)$-th coordinate equals to $\kappa_{\ell}(\mathbf{z},\mathbf{z}_{i}^{k}(\mathcal{X}_{\ell-1}^{(j)}))$. Similarly to Appendix D, denote $\tau_{\ell}^{f}(\mathbf{z}):=\sum_{k=0}^{K}\langle(Q_{\ell}^{k})^{\dagger}\mathbf{v}_{\ell}^{k}(\mathbf{z}),\sum_{s\in[F_{\ell-1}^{\star}]}\lambda_{\ell ks}w_{\ell ks}^{f}\mathbf{u}_{\ell ks}^{\top}\rangle$. For any matrix $\mathcal{Z}\in\mathbb{R}^{(K+1)\times F_{\ell-1}}$ whose $(k+1)^{\text{th}}$ row is given by $[\mathcal{Z}]_{k+1}:=\mathbf{z}_{k}$ for any $k\in[K]\cup\{0\}$, note that $\tau_{\ell}^{f}$ can be rephrased as $\tau_{\ell}^{f}(\mathbf{z}):=\sum_{k=0}^{K}\langle\mathbf{v}_{\ell}^{k}(\mathbf{z}),((Q_{\ell}^{k})^{\top})^{\dagger}\sum_{s\in[F_{\ell-1}^{\star}]}\lambda_{\ell ks}w_{\ell ks}^{f}\mathbf{u}_{\ell ks}^{\top}\rangle$. Thus, (22) implies that the Hilbert norm of the function $\tau_{\ell}^{f}$ satisfies:
$\|\tau_{\ell}^{f}\|_{\mathcal{H}_{\ell}}\leq\sum_{k=0}^{K}\Bigg{\|}(Q_{\ell}^{k})^{\top}((Q_{\ell}^{k})^{\top})^{\dagger}\sum_{s\in[F_{\ell-1}^{\star}]}\lambda_{\ell ks}w_{\ell ks}^{f}\mathbf{u}_{\ell ks}^{\top}\Bigg{\|}_{2}\leq\sum_{k=0}^{K}\Bigg{\|}\sum_{s\in[F_{\ell-1}^{\star}]}\lambda_{\ell ks}w_{\ell ks}^{f}\mathbf{u}_{\ell ks}^{\top}\Bigg{\|}_{2}$ (34)
$\leq\sum_{k=0}^{K}\sum_{s\in[F_{\ell-1}^{\star}]}|\lambda_{\ell ks}|\cdot|w_{\ell ks}^{f}|\cdot\|\mathbf{u}_{\ell ks}^{\top}\|_{2}$
$\leq\sum_{k=0}^{K}\sum_{s\in[F_{\ell-1}^{\star}]}|\lambda_{\ell ks}|=\sum_{k=0}^{K}\|\mathcal{A}_{\ell k}\|_{*}\leq C_{\sigma,\ell}(R_{\ell})(K+1)\sqrt{F_{\ell}F_{\ell-1}}$
where we used the triangle inequality and the fact that the singular vectors are unit vectors. Since $[\Psi^{\mathcal{A}_{\ell}}(\mathcal{X})]_{if}=\tau_{\ell}^{f}(\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1}))$, we obtained that $\mathcal{F}_{\mathcal{B}_{\ell}}\subset\mathcal{F}_{\text{cgcn}}$, as desired.
To conclude the proof, we next prove that any empirical risk minimizer $\Phi^{\mathcal{A}}$ in $\mathcal{F}_{\text{cgcn}}$ is also in $\mathcal{F}_{\mathcal{B}_{\ell}}$. By our proof in Appendix D, the value of the $f^{\text{th}}$ feature of node $i$ at layer $\ell$ is then given by $\tau_{\ell}^{f}(\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1}^{(j)}))$, where, as specified in (25), $\tau_{\ell}^{f}(\mathcal{Z})=\sum_{k=0}^{K}\langle(Q_{\ell}^{k})^{\dagger}\mathbf{v}_{\ell}^{k}(\mathbf{z}_{k}),(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f}\rangle$ is a reparameterization of the graph filters at layer $\ell$ for some vector of the coefficients $\mathbf{\beta}_{\ell k}^{f}\in\mathbb{R}^{nm}$ whose $(i,j)$-th coordinate is $\beta_{\ell k,(i,j)}^{f}$. That is, $(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f}$ now act as the filters under this reparameterization. In Appendix D, we have shown that the Hilbert norm of $\tau_{\ell}^{f}$ is then $\|\tau_{\ell}^{f}\|_{\mathcal{H}_{\ell}}=\|(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f}\|_{2}\leq C_{\sigma,\ell}(R_{\ell})$, where the last inequality is by (22). By (26), we further infer that the nuclear norm of the filters $(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f}$ in the new parameterization satisfies $\|(Q_{\ell}^{k})^{\top}\mathbf{\beta}_{\ell k}^{f}\|_{*}\leq C_{\sigma,\ell}(R_{\ell})\sqrt{F_{\ell}F_{\ell-1}(K+1)}\leq C_{\sigma,\ell}(R_{\ell})(K+1)\sqrt{F_{\ell}F_{\ell-1}}$, meaning that $\Phi^{\mathcal{A}}\in\mathcal{F}_{\mathcal{B}_{\ell}}$, as desired. ∎
Next, we prove that $\mathcal{F}_{L,cgcn}$ is not "too big" by upper bounding its Rademacher complexity. The Rademacher complexity is a key concept in empirical process theory and can be used to bound the generalization error in our empirical risk minimization problem. Readers should refer to [2] for a brief overview on Rademacher complexity. The Rademacher complexity of a function class =$\mathcal{F}=\{\phi:\mathcal{D}\rightarrow\mathbb{R}\}$ with respect to $m$ i.i.d samples $\{\mathcal{X}^{(j)}\}_{j=1}^{m}$ is given by:
$\mathcal{R}_{m}(\mathcal{F}):=\mathbb{E}_{\mathcal{X},\varepsilon}\Bigg{[}\sup_{\phi\in\mathcal{F}}\frac{1}{m}\sum_{j=1}^{m}\varepsilon_{j}\phi(\mathcal{X}^{(j)})\Bigg{]}$ (35)
where $\{\varepsilon_{j}\}_{j=1}^{m}$ are an i.i.d. sequence of uniform {±$\{\pm 1\}$-valued variables. Next, we provide the upper bound on the Rademacher complexity of the function class $\mathcal{F}_{L,cgcn}$ describing the graph filters. The lemma refers to the random kernel matrix $\mathcal{K}_{\ell}^{k}(\mathcal{X})\in\mathbb{R}^{n\times n}$ induced by a graph signal $\mathcal{X}\in\mathbb{R}^{n\times F}$ drawn randomly from the population, i.e., the $(i,i^{\prime})$-th entry of $\mathcal{K}_{\ell}^{k}(\mathcal{X})$ is $\kappa_{\ell}(\mathbf{z}_{i}^{k}(\mathcal{X}),\mathbf{z}_{i^{\prime}}^{k}(\mathcal{X}))$. Furthermore, we consider the expectation $\mathbb{E}[||\mathcal{K}_{\ell}^{k}(\mathcal{X})||_{2}]$ of this matrix’s spectral norm. While the previous lemmas consider the most general case where the number of features $F_{L}=G$ at the last layer may be any positive integer, observe that the following result focuses on the the binary classification case where $F_{L}=G=1$. As mentioned in Section 4.5, Lemma F.3 can be readily extended to the multi-class case by employing a standard one-versus-all reduction to the binary case.
Lemma F.3.
In the binary classification case where $F_{L}=G=1$, there exists a universal constant $c>0$ such that:
$\mathcal{R}_{m}(\mathcal{F}_{L,cgcn})\leq\frac{c\cdot C_{\sigma,\ell}(\mathcal{B}_{L})\sqrt{F_{\ell-1}(K+1)\cdot\log(n(K+1))\cdot\sum_{k=0}^{K}\mathbb{E}[||K_{L}^{k}(\mathcal{X})||_{2}]}}{\sqrt{m}}$ (36)
Proof.
We begin with some preliminaries that apply to the general case and then leverage them in order to obtain the result. For brevity, we denote $R_{\ell}=C_{\sigma,\ell}(\mathcal{B}_{\ell})\sqrt{F_{\ell}F_{\ell-1}}$. Consider some function $\Phi^{\mathcal{A}}=\phi^{{\mathcal{A}}_{L}}\circ\dots\circ\phi^{{\mathcal{A}}_{1}}\in\mathcal{F}_{\text{cgcn}}$. Without loss of generality, we assume that, each input input graph signal $\mathcal{X}^{(j)}$, the feature vector $[\mathcal{X}^{(j)}]_{i}=\mathbf{x}_{i}^{(j)}$ of each node $i$ resides in the unit $\ell_{2}$-ball (which can be obtain via normalization). Combined with $\kappa_{\ell}$’s being a continuous kernel with $\kappa_{\ell}(\mathbf{z},\mathbf{z})\leq 1$, Mercer’s theorem [44, Theorem 4.49] implies that there exists a feature map $\varphi_{\ell}:\mathbb{R}^{F_{\ell-1}}\rightarrow\ell^{2}(\mathbb{N})$ for which $\sum_{\xi=1}^{\infty}[\varphi_{\ell}(\mathbf{z})]_{\xi}[\varphi_{\ell}(\mathbf{z}^{\prime})]_{\xi}=\sum_{\xi=1}^{\infty}\varphi_{\ell\xi}(\mathbf{z})\varphi_{\ell\xi}(\mathbf{z}^{\prime})$ converges uniformly and absolutely to $\kappa_{\ell}(\mathbf{z},\mathbf{z}^{\prime})$. Accordingly, it holds that $\kappa_{\ell}(\mathbf{z},\mathbf{z}^{\prime})=\langle\varphi_{\ell}(\mathbf{z}),\varphi_{\ell}(\mathbf{z}^{\prime})\rangle$, which yields the following for each feature $f$ of any node $i$ at layer $\ell$ is $\tau_{\ell}^{f}(\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1}))$, where for any matrix $\mathcal{Z}\in\mathbb{R}^{(K+1)\times F_{\ell-1}}$ whose $(k+1)^{\text{th}}$ row is given by $[\mathcal{Z}]_{k+1}:=\mathbf{z}_{k}$ for any $k\in[K]\cup\{0\}$:
$\tau_{\ell}^{f}(\mathcal{Z}):=\sum_{k=0}^{K}\langle\varphi_{\ell}(\mathbf{z}_{k}),\bar{\mathbf{a}}^{f}_{\ell k}\rangle$ (37)
where $\bar{\mathbf{a}}^{f}_{\ell k}\in\ell^{2}(\mathbb{N})$ is a countable-dimensional vector, due to the fact that $\varphi$ itself is a countable sequence of functions. Particularly, the Hilbert norm of $\tau_{\ell}^{f}$ is $\|\tau_{\ell}^{f}\|_{\mathcal{H}_{\ell}}=\|\bar{\mathbf{a}}^{f}_{\ell k}\|_{2}$. Now, for any graph signal $\mathcal{X}$, let $\Theta_{i}(\mathcal{X})$ be the linear operator that maps any sequence $\theta\in\ell^{2}(\mathbb{N})$ to the vector $[\langle\varphi_{\ell}(\mathbf{z}_{i}^{0}(\mathcal{X})),\theta\rangle,\dots,\langle\varphi_{\ell}(\mathbf{z}_{i}^{K}(\mathcal{X})),\theta\rangle]^{\top}\in\mathbb{R}^{K+1}$. Thereby, recalling that $\Phi^{\mathcal{A}}=\phi^{{\mathcal{A}}_{L}}\circ\dots\circ\phi^{{\mathcal{A}}_{1}}$ where $\phi^{{\mathcal{A}}_{\ell}}\in\mathcal{F}_{\ell,cgcn}$ satisfies $[\phi^{{\mathcal{A}}_{\ell}}(\mathcal{X}_{\ell-1})]_{if}=\tau_{\ell}^{f}(\mathcal{Z}_{i,\ell}(\mathcal{X}_{\ell-1}))$, then the $f^{\text{th}}$ feature of node $i$ at layer $\ell$ can be rephrased as:
=${}_{if}=\mathbf{1}_{K+1}^{\top}\Theta_{i}(\mathcal{X}_{\ell-1})\bar{\mathbf{a}}^{f}_{\ell k}=\operatorname*{trace}\Big{(}\Theta_{i}(\mathcal{X}_{\ell-1})(\bar{\mathbf{a}}^{f}_{\ell k}\mathbf{1}_{K+1}^{\top})\Big{)}$ (38)
$\Rightarrow\phi^{{\mathcal{A}}_{\ell}}(\mathcal{X}_{\ell-1})=\Big{[}\operatorname*{trace}\Big{(}\Theta_{i}(\mathcal{X}_{\ell-1})(\bar{\mathbf{a}}^{f}_{\ell k}\mathbf{1}_{K+1}^{\top})\Big{)}\Big{]}_{i\in[n],f\in[F_{\ell}]}$
where $\mathbf{1}_{K+1}$ is the all-ones vector of dimension $K+1$. Note that the matrix $\bar{\mathcal{A}}_{\ell k}:=\bar{\mathbf{a}}^{f}_{\ell k}\mathbf{1}_{K+1}^{\top}$ satisfies the following:
$\|\bar{\mathcal{A}}_{\ell k}\|_{*}=\|\bar{\mathbf{a}}^{f}_{\ell k}\mathbf{1}_{K+1}^{\top}\|_{*}\leq\|\mathbf{1}_{K+1}^{\top}\|_{2}\cdot\|\bar{\mathbf{a}}^{f}_{\ell k}\|_{2}=\|\tau_{\ell}^{f}\|_{\mathcal{H}_{\ell}}\sqrt{K+1}\leq R_{\ell}\sqrt{K+1}$ (39)
Now, we obtain our upper bound for the binary classification case where $F_{L}=G=1$. Combining (38) with (39), we obtain that the Rademacher complexity of $\mathcal{F}_{L,cgcn}$ is upper bounded as follows:
$\mathcal{R}_{m}(\mathcal{F}_{L,cgcn})=\mathbb{E}\Bigg{[}\sup_{\phi^{{\mathcal{A}}_{L}}\in\mathcal{F}_{L,cgcn}}\frac{1}{m}\sum_{j=1}^{m}\varepsilon_{j}\Phi^{\mathcal{A}_{L}}(\mathcal{X}^{(j)})\Bigg{]}=$ (40)
$=\frac{1}{m}\mathbb{E}\Bigg{[}\sup_{\|\bar{\mathcal{A}}_{Lk}\|_{*}\leq R_{L}\sqrt{K+1}}\operatorname*{trace}\Big{(}\Big{(}\sum_{j=1}^{m}\varepsilon_{j}\Theta_{i}(\mathcal{X}_{L-1}^{(j)})\Big{)}\bar{\mathcal{A}}_{Lk})\Big{)}\Bigg{]}$
$\leq\frac{R_{L}\sqrt{K+1}}{m}\mathbb{E}\Bigg{[}\Big{\|}\sum_{j=1}^{m}\varepsilon_{j}\Theta_{i}(\mathcal{X}_{L-1}^{(j)})\Big{\|}_{2}\Bigg{]}$
where the last equality uses Hölder’s inequality while using the duality between the nuclear norm and the spectral norm. Next, note that $\sum_{j=1}^{m}\varepsilon_{j}\Theta_{i}(\mathcal{X}_{L-1}^{(j)})$ can be thought of as a matrix with $n$ rows and infinitely many columns. We denote its sub-matrix comprising of the first $r$ columns as $\Theta_{i}^{(r)}(\mathcal{X}_{L-1}^{(j)})$ and let $\Theta_{i}^{(-r)}(\mathcal{X}_{L-1}^{(j)})$ be the remaining sub-matrix. Therefore:
$\mathbb{E}\Bigg{[}\Big{\|}\sum_{j=1}^{m}\varepsilon_{j}\Theta_{i}(\mathcal{X}_{L-1}^{(j)})\Big{\|}_{2}\Bigg{]}$ (41a)
$\leq\mathbb{E}\Bigg{[}\Big{\|}\sum_{j=1}^{m}\varepsilon_{j}\Theta_{i}^{(r)}(\mathcal{X}_{L-1}^{(j)})\Big{\|}_{2}\Bigg{]}+\Bigg{(}\mathbb{E}\Bigg{[}\Big{\|}\sum_{j=1}^{m}\varepsilon_{j}\Theta_{i}^{(-r)}(\mathcal{X}_{L-1}^{(j)})\Big{\|}_{F}^{2}\Bigg{]}\Bigg{)}^{1/2}$ (41b)
$\leq\mathbb{E}\Bigg{[}\Big{\|}\sum_{j=1}^{m}\varepsilon_{j}\Theta_{i}^{(r)}(\mathcal{X}_{L-1}^{(j)})\Big{\|}_{2}\Bigg{]}+\Bigg{(}nm\mathbb{E}\Bigg{[}\sum_{\xi=r+1}^{\infty}(\varphi_{\ell\xi}(\mathbf{z}))^{2}\Bigg{]}\Bigg{)}^{1/2}$ (41c)
Since $\sum_{\xi=1}^{\infty}(\varphi_{\ell\xi}(\mathbf{z}))^{2}$ uniformly converges to $\kappa_{\ell}(\mathbf{z})$, the second term in (41c) converges to 00 as $r\rightarrow\infty$. Hence, it is sufficient to upper bound the first term in (41c) and take the limit $r\rightarrow\infty$. By Bernstein inequality due to [30, Theorem 2.1], whenever $\operatorname*{trace}(\Theta_{i}^{(r)}(\mathcal{X}_{L-1}^{(j)})\Theta_{i}^{(r)}(\mathcal{X}_{L-1}^{(j)})^{\top})\leq C_{1}$, there is a universal constant $c>0$ such that the expected spectral norm is upper bounded by:
$\mathbb{E}\Bigg{[}\Big{\|}\sum_{j=1}^{m}\varepsilon_{j}\Theta_{i}^{(r)}(\mathcal{X}_{L-1}^{(j)})\Big{\|}_{2}\Bigg{]}\leq c\sqrt{\log(mC_{1})}\mathbb{E}\Bigg{[}\Bigg{(}\sum_{j=1}^{m}\Big{\|}\Theta_{i}^{(r)}(\mathcal{X}_{L-1}^{(j)})\Theta_{i}^{(r)}(\mathcal{X}_{L-1}^{(j)})^{\top}\Big{\|}_{2}\Bigg{)}^{1/2}\Bigg{]}$ (42)
$\leq c\sqrt{\log(mC_{1})}(m\mathbb{E}[\|\Theta_{i}(\mathcal{X})\Theta_{i}(\mathcal{X})^{\top}\|_{2}])^{1/2}$
$\leq c\sqrt{\log(mC_{1})}\Bigg{(}m\sum_{k=0}^{K}\mathbb{E}[||K_{L}^{k}(\mathcal{X})||_{2}]\Bigg{)}^{1/2}$
Notice that, due to the uniform kernel expansion $\kappa_{\ell}(\mathbf{z},\mathbf{z}^{\prime})=\sum_{\xi=1}^{\infty}\varphi_{\ell\xi}(\mathbf{z})\varphi_{\ell\xi}(\mathbf{z}^{\prime})$, it holds that the trace norm is bounded as $\operatorname*{trace}(\Theta_{i}^{(r)}(\mathcal{X}_{L-1}^{(j)})\Theta_{i}^{(r)}(\mathcal{X}_{L-1}^{(j)})^{\top})\leq K+1$ since $\kappa_{\ell}(\mathbf{z},\mathbf{z}^{\prime})\leq 1$. Combining this with (42) and (40) implies the desired in (36). ∎
We are ready to prove the key lemma that will yield our main result. Recall that our current focus is on binary classification (i.e., $F_{L}=G=1$). While our prior results consider models with multiple hidden layers, our main result regards the case of a single hidden layer $L=1$. As mentioned in Section 4.5, in the case of multiple hidden layers, our result applies to each individual layer separately. Using Lemmas F.1F.3, we are capable of comparing the CGCN predictor obtained by Algorithm 2 against optimal model in the GCN class. Lemma F.2 yields that the predictor $\hat{\Psi}^{\hat{\mathcal{A}}_{L}}(\cdot)$ generated at the last layer is an empirical risk minimizer in the function class $\mathcal{F}_{\text{cgcn}}$. As such, by the theory of Rademacher complexity (see, e.g., [2]), we obtain that:
$\mathbb{E}_{\mathcal{X},\mathbf{y}}[J(\hat{\Psi}^{\hat{\mathcal{A}}_{L}}(\mathcal{X}),\mathbf{y})]\leq\inf_{\Phi\in\mathcal{F}_{L,cgcn}}\mathbb{E}_{\mathcal{X},\mathbf{y}}[J(\Phi(\mathcal{X}),\mathbf{y})]+2M\cdot\mathcal{R}_{m}(\mathcal{F}_{L,cgcn})+\frac{c}{\sqrt{m}}$ (43)
where $c$ is a universal constant a and we used the assumption that $J$ is $M$-Lipschitz. By Lemma F.1, we have that $\inf_{\Phi\in\mathcal{F}_{L,cgcn}}\mathbb{E}_{\mathcal{X},\mathbf{y}}[J(\Phi(\mathcal{X}),\mathbf{y})]\leq\inf_{\Phi\in\mathcal{F}_{L,gcn}}\mathbb{E}_{\mathcal{X},\mathbf{y}}[J(\Phi(\mathcal{X}),\mathbf{y})]$. Since $\mathcal{F}_{\text{cgcn}}=\mathcal{F}_{1,cgcn}$ due to $L=1$, this means that $\inf_{\Phi\in\mathcal{F}_{\text{cgcn}}}\mathbb{E}_{\mathcal{X},\mathbf{y}}[J(\Phi(\mathcal{X}),\mathbf{y})]\leq\inf_{\Phi\in\mathcal{F}_{\text{gcn}}}\mathbb{E}_{\mathcal{X},\mathbf{y}}[J(\Phi(\mathcal{X}),\mathbf{y})]$. Combined with (43), we obtain the desired.
Remark F.4.
The constant $C_{\sigma,L}(\mathcal{B}_{L})$ depends on the convergence rate of the polynomial expansion of the activation function $\sigma$ (See Appendix B). Algorithmically, knowing the activation function is not required to run Algorithm 2. Theoretically, however, the choice of $\sigma$ matters for Theorem 4.7 when comparing the predictor produced by Algorithm 2 against the best GCN, as it affects the representation power of such GCN. If a GCN with an activation function $\sigma$ performs well, CGCN will similarly generalize strongly.
Appendix G Additional Experimental Details
G.1 Dataset Descriptions
Table 2 provides a summary of the statistics and characteristics of datasets used in this paper. All datasets are accessed using the TUDataset class [31] in PyTorch Geometric [14] (MIT License).
Table 2: Statistics of the graph classification datasets used in our experiments.
Dataset #Graphs Avg. Nodes Avg. Edges #Classes
MUTAG 118 18 20 2
PTC-MR 344 26 51 2
PROTEINS 1,113 39 73 2
NCI1 4,110 30 32 2
Mutagenicity 4,337 30 31 2
G.2 Additional Implementation Details
Hardware and Software. All experiments could fit on one GPU at a time. Most experiments were run on a server with 4 NVIDIA L40S GPUs. We ran all of our experiments in Python, using the PyTorch framework [34] (license URL). We also make use of PyTorch Geometric (PyG) [14] (MIT License) for experiments with graph data.
Implementation Details. Particularly, all models were implemented using PyTorch Geometric (PyG) [14]. For GATv2 and GraphGPS, we use the official PyG implementations of their original papers, while for other baselines we use their PyG implementations, as standard in prior works (see, e.g., [32, 36]). Each neural architecture consists of $L\in\{2,6\}$ layers of the considered baseline, followed by a Set2Set global pooling [46] to aggregate the learned node-level representations to a graph-level embedding and a two-layer MLP for the final classification. The embedding dimension of each hidden layer is set to $64$, while non-convex models use a ReLU activation function between layers. For convex models, we use the Gaussian RBF kernel $\kappa(\mathbf{z},\mathbf{z}^{\prime}):=\exp(-\gamma\|\mathbf{z}-\mathbf{z}^{\prime}\|_{2}^{2})$ with $\gamma=0.2$, and compute an approximate kernel matrix via Nyström approximation [48] with dimension $P=64$ (Recall Section 4.4). As we want to evaluate the performance of GNN models in isolation, note that we do not use DropOut or batch normalization. Further, since we do not focus on the impact of attention heads on performance, all convex and non-convex versions of GAT, GATv2, GraphTrans and GraphGPS use their default configurations of attention heads.
Appendix H Additional Experimental Results
H.1 Heatmap of Classification Accuracy
Figure 2 provides a heatmap summarizing the accuracy of all models across the different datasets. It complements Table 1 by visually reinforcing that convex models generally outperform their non-convex counterparts across nearly all datasets, with consistent and major improvements even for models like GATv2 and GraphGPS. This reaffirms the strength and generality of our convexification framework. For example, on MUTAG, the convex CGIN model reaches 83.2% accuracy, outperforming the two-layer (60.0%) and six-layer (61.2%) GINs. Similarly, CGraphGPS achieves 83.7%, exceeding the best non-convex variant by over 30%. This trend holds across other datasets like PTC-MR, PROTEINS, and NCI1, where convex versions of CGraphSAGE, CGATv2, and CGraphGPS outperform non-convex baselines by 10–40%.
Figure 2: Heatmap of classification accuracy (%) across all models and datasets. Rows correspond to models, columns to datasets. The names of convex models are bolded to emphasize their performance.
H.2 Bar Plots
The bar plots in Figure 3 highlight the consistent performance advantage of convex models over their non-convex counterparts across multiple datasets in a more straightforward way than numerical data alone. They further underscore how convex architectures provide not just dataset-specific improvements, but a general trend of robustness and high accuracy, even in the absence of large training sets. This emphasizes that our convex models harness stable optimization landscapes, enabling reliable learning where non-convex models exhibit erratic performance due to training instabilities.
Figure 3: Accuracy comparison of two-layer convex GNNs (prefixed with ‘C’) against their standard two-layer and six-layer non-convex counterparts across five graph classification datasets. Bars represent the mean accuracy over four runs, and error bars denote the standard deviation. Datasets are ordered from left to right by increasing size.
H.3 Radar Plots
The radar plots in Figure 4 offer a holistic view of how convex models perform across diverse datasets. They reveal that convex variants not only perform strongly on small datasets like MUTAG and PTC-MR, but also remain competitive or superior on larger datasets such as NCI1 and Mutagenicity. This multi-dimensional consistency further challenges the notion that expressive representations require deep, non-convex architectures, and instead supports that convex shallow models, when carefully designed, can adapt well across a spectrum of tasks and data scales.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 4: Radar plots showing the performance of each two-layer convex model across five graph classification datasets. Each axis corresponds to a dataset, with higher values indicating better classification accuracy. These plots illustrate the consistency and robustness of convex models across diverse graph types.